### v1

# EllipSect: a surface brightness analysis tool for GALFIT 3

Abstract
EllipSect is a user-friendly Python analysis tool that generates publication-quality figures from the output of GALFIT 3, a popular galaxy surface brightness 2D-fitting algorithm. EllipSect allows the direct analysis of GALFIT 3 preliminary modeling, to guide optimal parameter selection. Furthermore, EllipSect generates non-parametric measurements of some useful parameters that are not provided by GALFIT 3, such as the effective radius resulting from multi-component fits, cusp radius, and the Petrosian radius. This paper provides examples and a quick guide for EllipSect.
Ellipsect es una herramienta de análisis en Python fácil de utilizar que genera figuras de calidad para publicación a partir de los archivos de salida de GALFIT 3, un algoritmo popular para el ajuste 2D del brillo superficial de galaxias. Ellipsect permite el análisis directo de modelos preliminares de GALFIT 3, para guiar la selección óptima de parámetros. Además, Ellipsect genera medidas no paramétricas de algunos parámetros útiles que no son proporcionados por GALFIT 3, como el radio efectivo resultante de ajustes de multicomponentes, radio de cúspide y radio de Petrosian. Este artículo proporciona ejemplos y una guía rápida para Ellipsect.
[…]
0.1 Introduction
GALFIT 3111For installation, usage, and general help visit https://users.obs.carnegiescience.edu/peng/work/galfit/galfit.html (Peng et al., 2002, 2010) is a widely used stand-alone-two-dimensional (2D) fitting code designed to model the surface brightness (SB) distribution of astronomical sources. GALFIT 3 applies 2D-parametric models to account for galaxies’ substructures and individual stars; this means more information is used than the 1D approach. Each subcomponent has its independent orientation; hence, the separation of components is more reliable. A common misconception suggests that 1D approaches work better in low signal-to-noise (S/N) regions or the presence of low-SB features; indeed, there isn’t any advantage from 1D over the 2D approaches due to the Poisson behavior of the signal [^2] See a more detailed discussion at https://users.obs.carnegiescience.edu/peng/work/galfit/TFAQ.html#misconception1 on 1D vs. 2D modeling.. The user provides the point spread function (PSF), and GALFIT 3 has built-in functions to model the PSF. GALFIT 3 convolves the PSF with the chosen profile to fit a given galaxy. This differs from most previous approaches that deconvolved the galaxy’s image, generated the isophotes, and then fit a 1D profile (Jedrzejewski, 1987).
GALFIT 3 built-in functions or profiles are the following [^3] See GALFIT 3 user’s manual, downloadable from https://users.obs.carnegiescience.edu/peng/work/galfit/galfit.html:
  1. 1.
    Sérsic (Sérsic, 1968): This profile is a generalization of the de Vaucouleurs profile and is typically used to fit bulges, bars, and elliptical galaxies by allowing the concentration index $n$ to vary:
    $I(r)=I_{e}\exp\left[-\kappa_{n}\left(\left(\frac{r}{r_{e}}\right)^{\frac{1}{n}}-1\right)\right],$ (1)
    where $I_{e}$ is the SB at the effective radius, $r_{e}$, which encloses half of the luminosity. The parameter $\kappa_{n}$ is related to $n$, so $r_{e}$ becomes the half-light radius; for $n\in\mathbb{R}$, $\kappa_{n}=2n-\frac{1}{3}+\frac{4}{405n}+\frac{46}{25515n^{2}}+O(n^{-3})$ gives an approximation (Ciotti & Bertin, 1999). See, for example, GALFIT 3 user’s manual and Graham & Driver (2005) for a full discussion on the properties of the Sérsic profile.
  2. 2.
    de Vaucouleurs (de Vaucouleurs, 1948): A particular case of the Sérsic profile (index $n=4$, $\kappa_{4}=-7.66925$), provides an excellent approximation to the SB profile of giant ellipticals. This profile has the following expression:
    $I(r)=I_{e}\exp\left[-7.66925\left(1-\left(\frac{r}{r_{e}}\right)^{1/4}\right)\right]$ (2)
  3. 3.
    Exponential (Patterson, 1940): This popular profile is used to fit the disks of spiral galaxies; it’s a particular case of the Sérsic profile with $n=1$:
    $I(r)=I_{0}\exp\left(-\frac{r}{r_{s}}\right),$ (3)
    where $I_{0}$ is the central SB and $r_{s}$ is called the scale length, which is related to the effective radius by $r_{e}=1.67835\,r_{s}$ for $n=1$. Freeman (1970) proposed a combination of a de Vaucouleurs bulge and an exponential disk to fit late-type galaxies.
  4. 4.
    Nuker (Lauer et al., 1995): This is a general profile that includes three power-law slopes, namely: $\gamma$ (inner), $\alpha$ (sharpness of the transition from inner to outer slopes), and $\beta$ (outer), they are free to vary and help to account for the large cores of very luminous early-type galaxies, including cD galaxies. This profile is expressed as follows:
[…]
    where $r_{b}$ is called the break radius, and it’s the place where the curvature of the profile reaches its maximum and $I_{b}$ is the SB at $r_{b}$.
    Recently, Baes (2020) showed that the Sérsic profile is a unique subgroup of the Nuker function; hence, the de Vaucouleurs profile, the exponential profile, and the Gaussian profile discussed above become specific cases of the Nuker profile, as well.
  5. 5.
    Edge-on disk (van der Kruit & Searle, 1981): This profile is used to fit the vertical structure of edge-on spiral galaxies. It results after considering an locally isothermal thin disk (van der Kruit & Searle, 1981):
    $I(r,z)=I_{0}\left(\frac{r}{r_{s}}\right)K_{1}\left(\frac{r}{r_{s}}\right)\operatorname{sech}^{2}\left(\frac{z}{z_{h}}\right),$ (5)
    where $I_{0}=I(0,0)$ is the central SB, $r_{s}$ is the scale length on the disk’s major axis, $z_{s}$ is the scale high on the $z$ axis perpendicular to the disk, and $K_{1}$ is the modified Bessel function of the second kind.
  6. 6.
    Ferrers [^4] Peng et al. (2010) called it modified Ferrer profile.: This model is named after the triaxial potential introduced by Norman Macleod Ferrers (Ferrers, 1877; Binney & Tremaine, 1987, pp. 61-62). The modified Ferrer profile programmed in GALFIT 3 is given by the following expression:
    $I(r)=I_{0}\left[1-\left(\frac{r}{r_{out}}\right)^{(2-\beta)}\right]^{\alpha},$ (6)
    which shows an almost flat core and an outer truncation, and whose sharpness of the truncation is set by $\alpha$, while the central slope is determined by $\beta$, $I_{0}$ is the central SB. This profile is only valid for $r\leq r_{out}$, i.e., $I(r\geq r_{out})\equiv 0$.
    This profile has been used to fit bars and lenses; it has eight free parameters. Notwithstanding, a Sérsic profile whose index is $n<0.5$ approximates the Ferrers profile; therefore, either profile can be used to model the SB of bars. Besides, the results are indistinguishable (e.g., Kim et al., 2015).
  7. 7.
    Empirical King (Elson, 1999): It is given as follows:
    $I(r)=I_{0}\left[1-\frac{1}{\left(1-\left(\frac{r_{t}}{r_{c}}\right)^{2}\right)^{\frac{1}{\alpha}}}\right]^{-\alpha}\left[\frac{1}{\left(1-\left(\frac{r}{r_{c}}\right)^{2}\right)^{\frac{1}{\alpha}}}-\frac{1}{\left(1-\left(\frac{r_{t}}{r_{c}}\right)^{2}\right)^{\frac{1}{\alpha}}}\right]^{\alpha},$ (7)
    where the core radius $r_{c}$, the truncation radius $r_{t}$, and the power law index $\alpha$ are free parameters; however, in the standard modified King profile, $\alpha=2$ is fixed. This profile fits the SB of globular clusters.
  8. 8.
    Gaussian profile is used to fit PSF profiles, the contribution of active galactic nuclei and bars; it’s a special case of the Sérsic profile when $n=0.5$. It is described by the following equation:
    $I(r)=I_{0}\exp\left(\frac{-r^{2}}{2\sigma^{2}}\right),$ (8)
    the scale is parametrized by the full with at half maximum ${\rm FWHM}=2\sqrt{2\ln{2}}\approx 2.355\,\sigma$, instead of $r_{e}$.
  9. 9.
    Moffat (Moffat, 1969) is a profile used to fit PSF profiles or the emission of active nuclei of galaxies. The following equation describes it:
    $I(r)=\frac{I_{0}}{\left(1+\left(\frac{r}{r_{d}}\right)^{2}\right)^{n}},$ (9)
    where $n$ is the concentration index; as for the Gaussian profile, the scale is given by $FWHM=2\left(2^{\frac{1}{n}}-1\right)^{\frac{1}{2}}\,r_{d}$. Nevertheless, the Gaussian profile is more widely used than the Moffat.
  10. 10.
    The user could provide the PSF as an image. However, care should be taken when selecting the stellar image’s centroid, intensity, and sampling (see GALFIT 3 user’s manual).
  11. 11.
    The background sky is modeled as a plane, with an optional gradient applied along the $x$ and $y$ directions.
  12. 12.
    Spiral modes are used to model the spiral arms or asymmetrical galaxies; see Peng et al. (2010) and GALFIT 3 user’s manual for more details.
With GALFIT 3, the user can apply multiple profiles to fit the following components: bulges, disks, bars, spiral arms, and asymmetries to one or several galaxies simultaneously. This gives a substantial advantage in crowded galaxy fields, such as compact groups or clusters of galaxies. Galaxies being extended objects are modeled better when the galaxy of interest is fitted together with its closely neighboring galaxies instead of masking. This has been shown through simulations (Häussler et al., 2007; Añorve, 2012).
GALFIT 3 applies a $\chi^{2}_{\nu}$ optimization scheme for parameter selection, including the background, generating parameter uncertainties, and produces a FITS (Flexible Image Transport System) image cube containing the image of the input object, the model image resulting after the fit, and the residual image. However, GALFIT 3 does not include a graphical tool to compare the input image with its resulting model image. Users commonly applied IRAF’s ellipse; nevertheless, this suite of programs became difficult to access after IRAF’s discontinuation in 2013; fortunately, IRAF’s ellipse is implemented in the Astropy library and new releases through. In any case, the main difficulty is that one needs to adequate GALFIT 3 ’s output format to the format admitted by external codes. However, this process could be time-consuming and prone to mistakes in handling.
Additionally, in crowded fields, neighboring galaxies near the target galaxy can be fitted simultaneously using GALFIT 3, resulting in more accurate outcomes. Häussler et al. (2007) demonstrated through simulations that masking galaxies in such fields leads to inaccurate fits for both the target galaxy and its neighbors. Improved fits were achieved when galaxies nearby were fitted simultaneously. However, this method is not feasible with IRAF’s ellipse task, which processes only one galaxy at a time.
Hence, we are introducing EllipSect, a new Python tool that generates SB profiles and extracts complementary information directly from GALFIT 3 ’s output. We provide an easy-to-use tool to help users select and improve their models by introducing additional components or removing redundant components. EllipSect ’s output includes data from the observed galaxy, and its model fits in the form of one-dimensional profiles. For a more straightforward assessment of the results, EllipSect can handle individual model components for a detailed analysis.
Additionally, EllipSect complements GALFIT 3 analysis by calculating integrated magnitudes, mean SB at $r_{e}$ ($\mu\right>_{e}$), the SB at $r_{e}$ ($\mu_{e}$), the radius at 90% of total light, Kron radius (Kron, 1980), Petrosian radius (Petrosian, 1976), bulge-to-total ratio, and the total effective radius ($r_{e}^{t}$) of a galaxy when multi-components are used.
This paper is organized as follows: §1 presents a general introduction to GALFIT 3 and the built-in profiles, §2 explains EllipSect ’s implemented algorithms. §3 shows applications and examples, §4 and §5 presents a discussion and conclusions, respectively. The Appendices show how to install and run EllipSect.
0.2 EllipSect
This section explains how EllipSect works. We have divided it into two parts: first, we explain the main algorithm, and then, we explain the photometric measurements. Throughout this article, we refer to the individual profiles that make a SB model the components, while the set of SB components is called the model.
0.2.1 The Algorithm
EllipSect was designed to be user-friendly for most users. It omits any direct interaction with the code or translation of GALFIT 3 ’s data format. The program is efficiently executed via the command line. It only requires the latest GALFIT 3 ’s model output files.
GALFIT 3’s output files consist of a data cube in FITS format, which includes the original image but chopped into the fit region, the model image, and residual image (i.e., galaxy image -- model image), a file named galfit.nn (where nn indicates the number of attempts run by GALFIT 3). This last file contains the parameters that fit each component of the model. Additionally, GALFIT 3 creates a log file including the model parameter values and their errors. EllipSect uses the GALFIT output file (galfit.nn) to parse the input magnitude zero point, plate scale, mask file name, output file, and model parameters.
The output GALFIT 3 file may contain information about the simultaneous fit of neighboring galaxies to the target galaxy. Using the distances among model components, EllipSect can discern which components belong to each galaxy.
EllipSect uses the same mask employed by GALFIT to remove the undesired pixels from the image. Once this is done, EllipSect reads the sky component from the GALFIT output file and subtracts it from the output galaxy and model images. The program divides the galaxy and its model into sectors equally spaced in angle around their centers. This is accomplished using the sectors_photometry routine from the mgefit555mgefit is a software to obtain Multi-Gaussian Expansion (MGE) fits galaxy images (Cappellari, 2002) Python library by Cappellari (2002). The routine also calculates the counts within each sector, extending outward testing the galaxy’s center until a specified minimum count level is reached. This is set by the MINLEVEL parameter of the routine. The default value is zero for this parameter. The ellipse is divided into sectors from the major axis to the minor axis, with the number of sectors set to a default value of $19$, resulting in sectors spanning an angle of $5^{\circ}$. Sectors around the galaxy are built using the position angle and ellipticity found in the model component of the output GALFIT file. If multiple components are fitted to the galaxy, the geometric information is taken from the last component in the model, as it is identified as the galaxy’s outermost component. This component is typically an exponential model, although this may vary depending on the specific galaxy model used. The user can change all these default parameters.
EllipSect generates publication-quality graphs. It produces two basic graphs in its simplest mode: one showing the averaged surface brightness along the major axis and another displaying the SB for various azimuthal angles across multiple plots (see examples in the Results section). Additionally, this mode generates a png image featuring the galaxy, model, and residual. EllipSect can also export a file containing the data points for each plot, allowing users to create custom visualizations with other plotting routines. Optionally, the SB profile for each plot component can include the galaxy (via the $--comp$ flag). This feature allows users to visualize the behavior of model components and identify irrelevant components or redundancies (see, for example, Figure 3).
Based on the mgefit manual, the SB is calculated according to the following formula666this does not include the extinction parameter :
$\mu_{I}=zpt+5\times\log(Pl_{sce})+2.5\times\log(t_{exp})-2.5\times\log(C_{0})+0.1,$ (10)
where $zpt$ is the magnitude zero point, $C_{0}$ are the $counts\times pixel^{-1}$, $Pl_{sce}$ is the plate scale, $t_{exp}$ is the exposition time, and $0.1$ is a correction for infinite aperture.
0.2.2 Output photometry
In addition to generating surface brightness plots, EllipSect calculates additional photometric variables from the models beyond those provided by GALFIT 3. These variables are stored in an output file and are derived after EllipSect extracts the count values for each sector. This section describes these variables and explains their computation.
EllipSect connects to NED through the command curl to download a xml file to extract the photometric corrections and compute the absolute magnitude according to the appropriate band. In the case of luminosity, it is calculated using the absolute magnitude of the Sun taken from Willmer (2018). This file extracts the distance modulus and the galactic extinction. It also obtains a conversion parameter used to compute the radius in Kpc units from angular units. SB dimming also corrects the mean SB at effective radius $\mu\right>_{e}$ and SB at the effective radius $\mu_{e}$. The only correction not applied is the k-correction since it needs the color information that is not always available 777see http://kcor.sai.msu.ru/. However, correction through galactic morphology can be used as well. It is important to note that corrections are applied only to the quantities in the output file; the single and multiple plots mentioned earlier are not adjusted for these corrections. However, users can apply a correction factor as a command-line argument or generate plots using the data from the EllipSect output files.
NED provides two distance modulus measurements: one derived from cosmological models and another obtained directly from the galaxy. By default, EllipSect uses the redshift-independent measurement to calculate the luminosity, if available. Otherwise, it uses the distance modulus based on cosmological models.
Most output photometric variables are calculated within the ellipse corresponding to the radius enclosing $90\%$ of the total light. Key variables include Bumpiness (Blakeslee et al., 2006), tidal parameter (Tal et al., 2009), reduced chi-squared ($\chi^{2}_{\nu}$) within this ellipse (as opposed to the value provided by GALFIT 3, which applies to the entire fitted image section), signal-to-noise ratio (S/N), residual sum of squares, number of degrees of freedom, number of free parameters, absolute magnitude, the bulge-to-total ratio, and luminosity.
The Bumpiness parameter quantifies the irregular structure of a galaxy by measuring deviations in its SB from a smooth SB model. Blakeslee et al. (2006) employed this parameter and the Sérsic index to classify galaxies into elliptical, S0, spiral, and irregular types.
The Bumpiness parameter is computed as follows:
$B=\frac{\left(\frac{1}{N}\sum_{r_{i,j}<2r_{e}}[I_{i,j}-S(x_{i},y_{j}|r_{e},n)]_{s}^{2}-\sigma_{s}^{2}\right)^{\frac{1}{2}}}{\frac{1}{N}\sum_{r_{i,j}<2r_{e}}S(x_{i},y_{j}|r_{e},n)},$ (11)
The sum is performed over two effective radii ($2r_{e}$), but in this case, it is applied within the ellipse region defined above. Here, $N$ represents the number of unmasked pixels included in the computation, $I_{i,j}$ is the galaxy intensity at pixel $(i,j)$, $S$ is the convolved surface brightness (SB) model, and $\sigma_{s}^{2}$ is the estimated photometric error, which is taken directly from the sigma image produced by GALFIT 3.
Another useful parameter is the Tidal (Tal et al., 2009) parameter is defined as:
$T_{gal}=\left[\left(\frac{I_{(x,y)}}{M_{(x,y)}}\right)-1\right],$ (12)
where $I_{(x,y)}$ are the galaxy pixel intensity values in (x, y) and $M_{(x,y)}$ represent the intensity values of the model.
This parameter measures the differences between the galaxy image and its corresponding model. Tal et al. (2009) used this to measure the tidal disturbances in elliptical galaxies and explored its relation with environmental properties, AGN activity, and color.
Additionally, suppose the galaxy has been fitted with multiple Sérsic components (or variants such as Gaussian and exponential models). In that case, the program gives $r_{e}^{t}$. While GALFIT 3 provides $r_{e}$ for each component, the program determines $r_{e}^{t}$ by summing the surface brightness model of all components, calculating half of the total flux, and identifying the corresponding radius using the bisection method.
The Akaike Information Criterion (AIC) (Akaike, 1974) and the Bayesian Information Criterion (BIC) (Schwarz, 1978) determines which model best fits the galaxy among a given set. Both criteria penalize adding new parameters by increasing the estimate of $\chi_{\nu}$. The model with the lowest AIC or BIC value is considered the best fit.
EllipSect computes both criteria and also calculates an alternative Bayesian Information Criterion (BIC) using the number of resolution elements instead of the number of pixels defined by $N_{res}=\frac{n_{pix}}{\pi(\text{FWHM})^{2}}$.
Finally, additional output parameters are computed for each model component, including the fraction of light (such as the bulge-to-total ratio for each component), $\mu\right>_{e}$, $\mu_{e}$, absolute magnitude, luminosity radius at 90% of the total light, Kron radius (Kron, 1980), and Petrosian radius (Petrosian, 1976).
Sky computation
An incorrect sky background estimation can lead to inaccurate surface brightness model parameters, especially the Sérsic index. An appropriate strategy is determining the sky background and fixing it during model fitting. This reduces the number of free parameters and prevents parameter degeneracy.
EllipSect provides two methods for computing the sky background: The Sky gradient method and the random box method. These methods are not used to estimate the photometric variables discussed above. Instead, they enable a comparison with the sky value computed by GALFIT 3, allowing the user to decide whether to refine the model using the newly derived sky value from EllipSect.
Sky gradient method. This method follows Barden et al. (2012). In this approach, EllipSect constructs concentric rings using the same angular parameters and ellipticity employed for the elliptical sectors described in 0.2.1. The user specifies the ring width, and EllipSect enlarges the rings’ major axes outward from the galaxy’s center. A mask file is used to exclude undesired pixels. Optionally, EllipSect can remove the top $80\%$ and bottom $20\%$ of sky values in each ring. Finally, it computes the gradient based on the mean sky values of each ring.
A negative gradient arises in the galaxy’s inner regions because the central counts exceed those in the outskirts. Beyond this zone, the gradient fluctuates randomly between positive and negative values (see Figure 5 in Barden et al. (2012)). EllipSect generates a fit image with sky rings to illustrate this gradient. It begins calculating the gradient from an initial radius and stops when it becomes positive for the second time, following the recommendation by Barden et al. (2012). At that point, EllipSect returns the sky value determined for this radius. Figure 1 shows an example of an elliptical ring used to compute the sky.
Figure 1: Example of sky computation using the gradient method. The ring around the bright cluster member of Abell 2029 indicates the radius at which the sky was computed for this galaxy. For visualization purposes, all pixels in that elliptical ring are assigned the same numerical value corresponding to the ellipse’s major-axis length.
Random box method
For this method, EllipSect calculates the sky using randomly placed boxes around the galaxy. To achieve this, it first creates an elliptical mask that covers the entire galaxy and excludes pixels in the original input mask file. By default, the geometry for this elliptical mask comes from the GALFIT 3 output file, although the user may modify it.
Next, the program positions boxes randomly between this elliptical mask and a specified maximum radius (often at the image boundary). It can also optionally remove the top 80% and bottom 20% of pixel values in these boxes. The final sky value is derived by computing the mean, median, and standard deviation of all pixels in the boxes. The default settings place 20 boxes, each measuring 20 pixels across. This method follows the approach described in Gao & Ho (2017).
0.3 Results: Examples of SB plots
EllipSect has been used to analyze galaxy images from 2MASS (Two Micron All Sky Survey) in which estimations of morphological and structural parameters have been obtained through photometric decompositions using GALFIT 3 (Ríos-López et al., 2021). Moreover, EllipSect was used in a study of dust nuclear structures of a sample of Active Galactic Nuclei (AGN) in the local Universe through the Hubble Space Telescope (HST) images, SB models, and radiative transfer simulations (Reyes-Amador, 2021). The code has also been used in reports, conference proceedings, professional papers, and undergraduate and graduate theses (Ríos-López, 2021, Valerdi M. et al. paper in progress; Ríos-López et al. paper in progress).
Here, we present examples of surface brightness profile plots for three images: Holm 15A with Kitt Peak National Observatory (KPNO), M61 with 2MASS, and NGC4261 with the Hubble Space Telescope (HST).
0.3.1 Kitt Peak: Holm 15A
This example shows the SB profile plots of Abell 85 brightest cluster galaxy (BCG), the cD galaxy Holm 15A. Here, GALFIT 3 was used to fit a multi-gaussian component to model its SB; another analysis using the Nuker profiles has already been published (López-Cruz et al., 2014). This frame was taken with KPNO 0.9m Telescope, part of the catalog of LOCOS (López-Cruz, 1997). Details on the selection criteria, observations, and data reduction can be found in López-Cruz (1997), López-Cruz et al. (2004), and López-Cruz et al. (2025, in preparation). In Figure 2, we show the cube image generated by EllipSect for this galaxy.
Figure 2: Example of cube image generated by EllipSect. The left panel shows the galaxy, the middle panel is the model, and the right is the residual. The user can change the color palette, brightness, and contrast.
In Figure 3, we present the single-plot and multiplot of the surface brightness for the brightest cluster member of Abell 85. This galaxy was modeled using seven Gaussians to capture its full SB profile. The multiplot displays the error, while the single-plot shows the percent error. These plots were generated automatically by EllipSect, which computes the SB profiles of the galaxy, model, and individual components similar to IRAF’s ellipse task.
Figure 3: Example of EllipSect output for the elliptical galaxy Holm 15a and its model. This was fitted using seven Gaussian components. In both panels, red represents the galaxy and blue represents the GALFIT 3 model. The color for each Gaussian component is shown in the top right inset box. These colors represent the same for both panels. Left panel: average SB vs. radius along the major axis. The residual percentage is shown in the bottom left plot. The model errors depend on the average over the ellipse for the model in that radius. Right panel: Multiple plots of galaxy SB and model at different angles from the major axis (the major axis is the one with $0\deg$). The error is shown on the right side of the multiplot. The angle in red at the top right indicates the angle measured from the Y-axis of the image. On the other hand, the subtle grey angle at the bottom left indicates the angle measured from the major axis starting from the top with the major axis and the bottom plot with the minor axis
0.3.2 2MASS: M61
In Figure 4, we show the galaxy, model, and residual for galaxy M 61 in the Ks band of 2MASS.
Figure 4: Example of cube image generated by EllipSect for M 61 galaxy. Details for this image are the same as the caption of Figure 2.
Figure 5 shows the multiplot of the fitting of M 61. This model consists of a series of Sérsic and Exponential components. As can be seen, deviations from the model can be appreciated in the region where the model fails to reproduce the SB model of the galaxy.
Figure 5: Example of EllipSect output for M 61 galaxy and its model. The description of the image is the same as in the right panel of Figure 3
0.3.3 HST: NGC4261
In the last example, we show an image from HST; this is NGC4261 and in Figure 6 is shown the galaxy, model, and residual.
Figure 6: Example of cube image generated by EllipSect for NGC4261 galaxy. Details for this image are the same as the caption of Figure 2.
Figure 7 shows the galaxy NGC4261 taken with the HST using filter F547M. This model used a single Sersic component to fit the galaxy’s SB.
Figure 7: Example of EllipSect output for NGC4261 galaxy and its model. The description of the image is the same as the right panel of Figure 3.
0.4 Discussion
EllipSect was developed as an extension to GALFIT 3 to generate SB profile plots and offer additional photometric information beyond GALFIT 3. Its primary purpose is to test the reliability of GALFIT 3 ’s model at various radii and angles for a given galaxy. In addition, EllipSect allows users to include individual components of the GALFIT 3 models in the plots, enabling verification that these components behave as expected. This feature streamlines the analysis process by eliminating the need to switch between different program formats, making it practical to quickly confirm that the GALFIT model is appropriate for the galaxy. EllipSect aids GALFIT 3 users with a readily accessible and easy-to-use tool for generating accurate SB measurements and publication-quality plots.
Using the sectors_photometry routine, EllipSect creates evenly spaced angular sectors to compute surface brightness at various radii. It divides an ellipse, centered on the galaxy, into multiple sectors according to the geometric information from the GALFIT 3 output file. By default, EllipSect derives the angular orientation and ellipticity from the last component of the GALFIT 3 output file (see 0.6.2).
In the conventional approach to generating SB profiles, one typically uses IRAF’s ellipse. This routine fits an ellipse to each isophote at every radius, from the central region to the galaxy’s outskirts. As the radius changes, the ellipticity and position angle of the ellipse may also vary.
One disadvantage of EllipSect is that it maintains a fixed ellipticity and angular position values throughout the entire galaxy, which may hinder studying how these parameters change with radius. However, this fixed parameter approach is advantageous when dealing with galaxies simultaneously fitted by GALFIT 3. In such cases, masking or removing the galaxy is insufficient, as the galaxies are blended in the line of sight, and it is challenging to distinguish where one galaxy ends and the other begins (Haussler et al. 2007). In addition, EllipSect avoids the need for ”double” fitting, i.e., two-dimensional and one-dimensional fitting, which IRAF’s ellipse requires. One-dimensional routines may not be capable of simultaneous fitting of isophotes for two neighboring galaxies, making simultaneous fitting difficult or impossible. While EllipSect is not a complete substitute for IRAF’s ellipse, it offers distinct advantages depending on the scientific question being addressed.
Despite the widespread use of IRAF’s ellipse, alternative techniques such as EllipSect offer unique advantages for SB analysis, which may prove to be beneficial for astronomers seeking to enhance their understanding of galactic structures.
An important output parameter from EllipSect is $r_{e}^{t}$, the effective radius of the entire galaxy, which can be used to parameterize metallicity gradients and other relevant information (e.g., Garduño et al., 2023). If the galaxy has only a single Sérsic component, $r_{e}$ from that component can represent the whole galaxy. However, no single component’s $r_{e}$ can accurately describe the entire galaxy when multiple components are present, because galaxies do not perfectly follow the surface models.
To resolve this, EllipSect employs a numerical method that adds together the models of all components and then locates the radius where the total luminosity reaches half of its overall value ($r_{e}^{t}$). This procedure can also be applied to determine the radius corresponding to any percentage of the total light.
Determining which aperture to use for a galaxy’s photometric variables can be somewhat arbitrary. Different photometric quantities require a suitable radius for proper evaluation, and standardizing this across various galaxies is difficult. In the case of GALFIT 3, the reduced chi-squared ($\chi^{2}_{\nu}$) is computed using all pixels within the fitting region. However, changing the size of this region can alter $\chi^{2}_{\nu}$ and may include pixels from other objects, introducing additional noise. Consequently, the $\chi^{2}_{\nu}$ reflects the fit quality for every object within that region.
In contrast, EllipSect uses the radius containing 90% of the total light to compute various photometric variables. This radius is found by applying the same procedure used to determine the effective radius but at 90% of the total light. Focusing on this region ensures that only pixels belonging to the galaxy are included, minimizing contamination from external objects. Within this radius, EllipSect calculates several photometric variables, such as the tidal parameter, Bumpiness, AIC, BIC, and a local $\chi^{2}_{\nu}$.
EllipSect provides corrections for galaxy extinction and surface brightness (SB) dimming, but does not perform k-corrections at this time. Users can still apply morphological-based corrections to the galaxy, although this may be imprecise for cluster galaxies whose colors can differ from those of field galaxies. In such situations, manual correction is recommended. Note that k-corrections can be generated using online tools such as http://kcor.sai.msu.ru/.
EllipSect provides a helpful tool for wrapping codes aimed at automating the GALFIT 3 fitting process. It can efficiently compute the necessary photometric variables required by these codes, thereby allowing for rapid analysis of model outputs for multiple galaxies, which would otherwise be time-consuming. In fact, EllipSect has been specifically designed to be used in conjunction with the Driver for GALFIT on Cluster Galaxies (DGCG) code (Añorve et al., in preparation), which automates GALFIT 3 fitting for galaxies in cluster images.
0.5 Conclusions and Future Work
The paper introduces EllipSect, a Python tool for analyzing GALFIT 3 output files. It generates surface brightness profiles for galaxies, models, and individual components, offering single-plot and multiplot visualizations. Additionally, EllipSect computes various photometric parameters, including light fractions for each component, $r_{e}^{t}$, $\mu\right>_{e}$, $\mu_{e}$, absolute magnitudes, the luminosity radius $90\%$, the effective radius of the galaxy, the Kron radius, and the Petrosian radius. It also calculates tidal and Bumpiness parameters, signal-to-noise ratios, and local $\chi^{2}_{\nu}$ values to assess differences between models and observed galaxies.
EllipSect offers two methods for computing the sky background and enables users to distinguish among different models for galaxies by calculating the Akaike and Bayesian information criteria. The code can be easily adapted to other GALFIT 3 wrapping codes and will be integrated with DGCG in future work. This project aims to apply EllipSect to analyze data images containing hundreds of galaxies, such as galaxy clusters.
EllipSect is a robust tool for analyzing GALFIT 3 output and extracting extensive photometric information for galaxies. Its flexibility and versatility make it an essential tool for studying galaxy properties and morphology. Future development will focus on expanding its capacity to handle larger datasets and integrate it with other GALFIT 3 wrapping codes.
{appendices}
0.6 How to install and run EllipSect
0.6.1 Installing EllipSect
The code is written for Python 3. It can be downloaded from the site:
https://github.com/canorve/EllipSect
Quick installation
The quickest way to install EllipSect is just by typing this into the command line:
1pip install ellipse
Step-by-step installation
The following libraries are required to install EllipSect (v3.4.4) along with Python >=>=> = v3.8:
  • numpy >=>=> = v1.20.3
  • astropy >=>=> = v5.1
  • scipy >=>=> = v1.5.2
  • matplotlib >=>=> = v3.5.2
  • mgefit >=>=> = v5.0.13
  • sphinx >=>=> = v3.2.1
After installing the libraries, download the latest release from the link shown above, and install it via:
1cd EllipSect
2pip install .
or
1cd EllipSect
2python setup.py install
To check that everything is working properly, run the automated tests:
1python -m pytest
Although GALFIT is not strictly required, it will require to create the model components and sigma image. Ensure you can call GALFIT from the command line. Otherwise, the automated tests will fail.
0.6.2 Running EllipSect
Easy run: Once installed, run the command EllipSect in the same directory that you run GALFIT. The easiest way to run the program is:
1EllipSect GALFIT.01
Note: EllipSect takes the GALFIT output files (GALFIT.XX) where XX is the output file number you want to plot.
More options: As explained in the paper, EllipSect uses an ellipsed-based grid to compute the SB. The information of this ellipse is obtained from the GALFIT output file, but if the user wants to provide their ellipse parameters, this is done via:
1EllipSect GALFIT.01 -q 0.5 -pa 80
In this last example, an ellipse with a $0.5$ axis ratio and position angular of $80$ (measured from Y-axis) was used to create the plots.
In case your model contains two or more components, the comp (or cp) option will plot each component (for an example of this check figure 3):
1EllipSect GALFIT.01 -cp
There are many options to change the format of the plots such as: logx, ranx, rany, pix, grid, dpi, galax, but if the user prefers to make their plots the option sbout will create a file (or files) containing the points of the the SB of the galaxy and model at each radius.
To compute the photometry of the galaxy and model such as snr, BIC, AIC, absolute Magnitude, Luminosity, Bumpiness, Tidal and other photometric variables the option phot creates a file with the computed variables:
1 EllipSect GALFIT.01 --phot
EllipSect has the option to compute the sky to compare with or provide to GALFIT. As explained above, the program uses two methods to compute it: The gradient method and the random box method. To compute the gradient method:
1 EllipSect GALFIT.01 -gsky
And to compute using the random box method:
1 EllipSect GALFIT.01 -rsky
The ring’s width and box’s size can be modified from their default options.
For more options and to see how to use them, type:
1EllipSect --help
Running EllipSect as a script
If you want to use EllipSect inside your Python script, you can call it in the following way:
1 from ellipsect.inout.read import ArgParsing
2 from ellipsect.sectors.sect import SectorsGalfit
3
4 args=[’GALFIT.01’,’--logx’, ’--phot’,’--noplot’]
5 parser_args = ArgParsing(args)
6 photapi = SectorsGalfit(parser_args)
7
8 print("Akaike Criterion: ", photapi.AICrit)
9 print("Bulge to Total: ", photapi.BulgeToTotal)
Listing 1: Example Python Code with EllipSect
A manual and more details can be found in https://github.com/canorve/EllipSect

### v2

@@
-EllipSect: a surface brightness analysis tool for GALFIT 3
-Abstract
-EllipSect is a user-friendly Python analysis tool that generates publication-quality figures from the output of GALFIT 3, a popular galaxy surface brightness 2D-fitting algorithm.
-EllipSect allows the direct analysis of GALFIT 3 preliminary modeling, to guide optimal parameter selection.
-Furthermore, EllipSect generates non-parametric measurements of some useful parameters that are not provided by GALFIT 3, such as the effective radius resulting from multi-component fits, cusp radius, and the Petrosian radius.
-This paper provides examples and a quick guide for EllipSect.
-Ellipsect es una herramienta de análisis en Python fácil de utilizar que genera figuras de calidad para publicación a partir de los archivos de salida de GALFIT 3, un algoritmo popular para el ajuste 2D del brillo superficial de galaxias.
-Ellipsect permite el análisis directo de modelos preliminares de GALFIT 3, para guiar la selección óptima de parámetros.
-Además, Ellipsect genera medidas no paramétricas de algunos parámetros útiles que no son proporcionados por GALFIT 3, como el radio efectivo resultante de ajustes de multicomponentes, radio de cúspide y radio de Petrosian.
-Este artículo proporciona ejemplos y una guía rápida para Ellipsect.
+arXiv:2501.10079v2[astro-ph.IM] 02 Mar 2026
+arXiv:2501.10079v2[astro-ph.IM] 02 Mar 2026
 […]
 […]
-0.1 Introduction
-GALFIT 3111For installation, usage, and general help visit https://users.obs.carnegiescience.edu/peng/work/galfit/galfit.html (Peng et al., 2002, 2010) is a widely used stand-alone-two-dimensional (2D) fitting code designed to model the surface brightness (SB) distribution of astronomical sources.
-GALFIT 3 applies 2D-parametric models to account for galaxies’ substructures and individual stars; this means more information is used than the 1D approach.
-Each subcomponent has its independent orientation; hence, the separation of components is more reliable.
-A common misconception suggests that 1D approaches work better in low signal-to-noise (S/N) regions or the presence of low-SB features; indeed, there isn’t any advantage from 1D over the 2D approaches due to the Poisson behavior of the signal [^2] See a more detailed discussion at https://users.obs.carnegiescience.edu/peng/work/galfit/TFAQ.html#misconception1 on 1D vs.
-2D modeling..
-The user provides the point spread function (PSF), and GALFIT 3 has built-in functions to model the PSF.
+EllipSect: A Surface Brightness Analysis Tool for GALFIT 3
+C.
+Añorve Facultad de Ciencias de la Tierra y el Espacio, Universidad Autónoma de Sinaloa, Blvd.
+de la Américas y Av.
+Universidad s/n, Ciudad Universitaria, Culiacán, Sinaloa, 80013, México.
+O.
+U.
+Reyes-Amador Instituto de Radioastronomía y Astrofísica, UNAM, Campus Morelia, AP 3-72, C.P.
+58089, México.
+E.
+Ríos-López Instituto Nacional de Astrofísica, Óptica y Electrónica (INAOE), Luis Enrique Erro 1, Apartado Postal 51 y 216, C.P.
+72840 Puebla, México.
+Instituto de Astrofísica de Canarias (IAC), Vía Láctea s/n, 38205, La Laguna, Tenerife, España.
+Universidad Pedagógica del Estado de Sinaloa, Castiza s/n, 80027, Culiacán, Sinaloa, México.
+D.
+de Ramón Tadeo Facultad de Ciencias de la Computación, Benemérita Universidad Autónoma de Puebla (FCC-BUAP), Av.
+San Claudio y 14 Sur, Ciudad Universitaria, C.P.
+72570, Puebla, México.
+O.
+López-Cruz Instituto Nacional de Astrofísica, Óptica y Electrónica (INAOE), Luis Enrique Erro 1, Apartado Postal 51 y 216, C.P.
+72840 Puebla, México.
+canorve@uas.edu.mx
+keywords:
+galaxies: photometry, galaxies: fundamental parameters, methods: data analysis
+doi:
+https://www.astroscu.unam.mx/rmaa/RMxAA..ELLIPSECT
+EllipSect es una herramienta de análisis y medición, implementada en Python, que opera sobre los datos de imagen junto con la salida del código ampliamente utilizado de ajuste bidimensional de brillo superficial GALFIT 3.
+Genera figuras de calidad publicable y productos de datos exportables para permitir una evaluación cuantitativa de los modelos de GALFIT 3 y de sus componentes individuales.
+Además, EllipSect calcula mediciones no paramétricas que no son proporcionadas por GALFIT 3, incluyendo el radio efectivo total resultante de ajustes multicomponente, el radio de cúspide y el radio de Petrosian.
+Este artículo presenta ejemplos y una guía rápida de uso de EllipSect.
+1 Introduction
+GALFIT 3111For installation, usage, and general help, visit https://users.obs.carnegiescience.edu/peng/work/galfit/galfit.html (peng02; peng10) is a widely used stand-alone two-dimensional (2D) fitting code designed to model the surface brightness (SB) distribution of astronomical sources by generating 2D-parametric models that account for galaxies’ substructures and individual stars, among others.
+With GALFIT 3, the user can apply multiple profiles to fit the following components: bulges, disks, bars, spiral arms, and asymmetries to one or several galaxies simultaneously.
+Furthermore, when dealing with multiple components, the orientation of each component is independent of the others; hence, component-based separation is more reliable.
+This property gives a substantial advantage in crowded galaxy fields, such as compact groups or clusters of galaxies.
+Indeed, our experience shows that masking does not adequately account for light from nearby or overlapping galaxies; instead, fitting the galaxy of interest along with its neighboring galaxies yields a better result.
+Indeed, haussler07; anorve12 tested this recommendation using simulated data and found that fitting neighboring galaxies simultaneously accounts for galaxies’ extended light more precisely.
+Another feature of GALFIT 3 is that a user-supplied point spread function (PSF) image is necessary, or the user may utilize the built-in functions provided by GALFIT 3 to model the PSF.
 GALFIT 3 convolves the PSF with the chosen profile to fit a given galaxy.
-This differs from most previous approaches that deconvolved the galaxy’s image, generated the isophotes, and then fit a 1D profile (Jedrzejewski, 1987).
-GALFIT 3 built-in functions or profiles are the following [^3] See GALFIT 3 user’s manual, downloadable from https://users.obs.carnegiescience.edu/peng/work/galfit/galfit.html:
+This differs from some previous methods, which relied on deconvolving the image (for example lucy74; richardson72) before generating isophotes (e.g., jedrzejewski87) and fitting the SB profile.
+GALFIT 3 applies a $\chi^{2}_{\nu}$ optimization scheme for parameter selection, including the background, generating parameter uncertainties, and producing a FITS (Flexible Image Transport System) image cube containing the image of the input object, the model image resulting after the fit, and the residual image.
+However, GALFIT 3 does not include a graphical tool to compare the input image with its resulting model image.
+Users commonly applied one-dimensional (1D) IRAF’s ellipse and other plotting programs; nevertheless, this suite of programs became difficult to access after IRAF’s discontinuation in 2013.
+Fortunately, IRAF’s ellipse is implemented in the Astropy library and is available in new releases.
+In any case, the main difficulty is that one must convert GALFIT 3’s output format to a format accepted by external codes.
+Nevertheless, this process could be time-consuming and error-prone.
+However, this method is not feasible with IRAF’s ellipse task, which processes only one galaxy at a time.
+Also, a common misconception suggests that 1D approaches work better in low signal-to-noise (S/N) regions or the presence of low-SB features; however, there is no advantage from 1D over the 2D approaches due to the Poissonian nature of the signal [^2] See a more detailed discussion at https://users.obs.carnegiescience.edu/peng/work/galfit/TFAQ.html#misconception1 on 1D vs.
+2D modeling (see also press2007; peng10).
+Hence, we are introducing EllipSect, a new algorithm that generates SB profiles and extracts complementary information directly from GALFIT 3 ’s output, with the current implementation coded in Python.
+The current manuscript focuses on EllipSect as a post-fit analysis tool that produces quantitative measurements, publication-quality graphics, and exportable data products from the observed galaxy and GALFIT 3 models.
+These products can also support iterative model refinement (e.g., adding or removing components) when needed.
+EllipSect ’s output includes data from the observed galaxy, and its model fits in the form of one-dimensional profiles.
+For a more straightforward assessment of the results, EllipSect can handle individual model components for a detailed analysis.
+Additionally, EllipSect provides non-parametric measurements derived from the data and the GALFIT 3 model/output, such as integrated magnitudes, mean SB at $r_{e}$ ($\left<\mu\right>_{e}$), the SB at $r_{e}$ ($\mu_{e}$), the radius at 90% of total light, Kron radius (kron80), Petrosian radius (petrosian76), bulge-to-total ratio, and the total effective radius ($r_{e}^{t}$) of a galaxy with multiple components fits.
+To make this paper self-contained, Appendix A provides a brief history, applications, and analytical profiles implemented in GALFIT 3.
+We intended to order them logically, from the more general instance through their particular cases, to avoid the impression that the suite of profiles included in GALFIT 3 is merely a simple list of unrelated profiles.
+This paper contains the following sections: §1 presents a general introduction to GALFIT 3 and the built-in profiles; §2 explains EllipSect’s implemented algorithms.
+§3 shows applications and examples, §4 and §5 presents a discussion and conclusions, respectively.
+The appendices show the build-up of the GALFIT 3 functions and explain how to install and run EllipSect.
+2 EllipSect
+This section explains how EllipSect works.
+We have divided it into two parts: first, we describe the main algorithm, and then, we explain the photometric measurements.
+Throughout this article, we refer to the individual profiles that comprise a SB model as the components.
+In contrast, the set of SB components constitutes the model.
+2.1 The Algorithm
+EllipSect was designed to be user-friendly for most users.
+It omits any direct interaction with the code or translation of GALFIT 3 ’s data format.
+EllipSect is executed from the command line.
+It requires only the latest GALFIT 3 model output files.
+GALFIT 3’s output files consist of a data cube in FITS format, which includes the original image (chopped into the fit region), the model image, and the residual image (i.e., galaxy image - model image), as well as a file named galfit.nn (where nn indicates the number of attempts run by GALFIT 3).
+This last file contains the parameters that fit each component of the model.
+Additionally, GALFIT 3 creates a log file that includes the model parameter values and their associated errors.
+EllipSect uses the GALFIT output file (galfit.nn) to parse the input magnitude zero point, plate scale, mask file name, output file, and model parameters.
+The output GALFIT 3 file may include information about the simultaneous fits of neighboring galaxies to the target galaxy.
+Using the distances among the model components, EllipSect can identify the elements associated with each galaxy.
+EllipSect uses the same mask employed by GALFIT to remove the undesired pixels from the image.
+Once this is done, EllipSect reads the sky component from the GALFIT output file and subtracts it from the output galaxy and model images.
+The program divides the galaxy and its model into sectors equally spaced in angle around their centers, using the sectors_photometry routine from the mgefit333mgefit is a software to obtain Multi-Gaussian Expansion (MGE) fits of galaxy images (cappellari02) Python library by cappellari02.
+With this routine, EllipSect employs these sectors to compute the surface brightness at multiple radii.
+It divides an ellipse, centered on the galaxy, into multiple sectors according to the geometric information from the GALFIT 3 output file until reaching a minimum count level, set by the MINLEVEL parameter, and its preselected value is zero.
+EllipSect takes the angular orientation and ellipticity from the last component of the GALFIT 3 output file (see 7.2) as default values.
+This component is typically an exponential model, although this may vary with the specific galaxy model employed.
+The user can change all default parameters.
+EllipSect generates publication-quality graphs.
+It produces two basic graphs in its simplest mode: one showing the averaged surface brightness along the major axis and another displaying the SB for various azimuthal angles across multiple plots (see examples in the Results section).
+Additionally, this mode generates a png image featuring the galaxy, model, and residual.
+EllipSect can also export a file containing the data points for each plot, allowing users to create custom visualizations with other plotting routines.
+Optionally, the SB profile for each plot component can include the galaxy (via the --comp flag).
+This feature allows users to visualize the behavior of model components and identify irrelevant components or redundancies (see, for example, Figure 3).
+According to the mgefit manual, the SB results from the following expression [^4] This does not include the extinction parameter:
+$\mu_{I}=zpt+5\times\log(Pl_{sce})+2.5\times\log(t_{exp})-2.5\times\log(C_{0})+0.1,$ (1)
+where zptzpt is the magnitude zero point, $C_{0}$ are the $counts\times pixel^{-1}$, $Pl_{sce}$ is the plate scale, $t_{exp}$ is the exposition time, and $0.1$ is a correction for infinite aperture.
+2.2 Output photometry
+In addition to generating surface brightness plots, EllipSect calculates additional photometric variables from the models beyond those provided by GALFIT 3.
+These variables are stored in an output file and are computed after EllipSect extracts the counts for each sector.
+This section describes these variables and explains how they are calculated.
+EllipSect connects to NED through the command curl to download an xml file to extract the photometric corrections and compute the absolute magnitude according to the appropriate band.
+To transform magnitudes to luminosities, and vice-versa, the absolute magnitude of the Sun from willmer18 is assumed.
+This file extracts the distance modulus and the galactic extinction (schlafly11).
+It also retrieves a conversion parameter used to convert angular units to Kpc.
+SB dimming also corrects the mean SB at effective radius $\left<\mu\right>_{e}$ and SB at the effective radius $\mu_{e}$.
+The only correction not applied is the k-correction, as it needs the color information that is not always available (e.g., see chilingarian12).
+It is important to note that corrections are applied only to the quantities in the output file, while the single and multiple plots mentioned earlier do not include these corrections.
+However, users can apply a correction factor either as a command-line argument or by generating plots from the EllipSect output files.
+NED provides two distance-modulus measurements: one derived from cosmological models and the other obtained directly from the galaxy.
+By default, EllipSect uses the redshift-independent measurement to calculate the luminosity, if available.
+Otherwise, it uses the distance modulus based on cosmological models.
+The majority of photometric parameters are calculated within an ellipse that encloses 90% of the total light.
+Some of those associated photometric parameters are the Bumpiness (blakeslee06), the Tidal parameter (tal09), the reduced chi-squared ($\chi^{2}_{\nu}$) within this ellipse (as opposed to the value provided by GALFIT 3, which applies to the entire fitted image section), the signal-to-noise ratio (S/N), the residual sum of squares, the number of degrees of freedom, the number of free parameters, the absolute magnitude, the bulge-to-total ratio, and the luminosity.
+The Bumpiness parameter quantifies the irregular structure of a galaxy by measuring deviations in its SB from a smooth SB model.
+blakeslee06 employed this parameter and the Sérsic index to classify galaxies into elliptical, S0, spiral, and irregular types.
+The calculation of the Bumpiness $B$ results from the following formula:
+$B=\frac{\left(\frac{1}{N}\sum_{r_{i,j}<2r_{e}}[I_{i,j}-S(x_{i},y_{j}|r_{e},n)]_{s}^{2}-\sigma_{s}^{2}\right)^{\frac{1}{2}}}{\frac{1}{N}\sum_{r_{i,j}<2r_{e}}S(x_{i},y_{j}|r_{e},n)},$ (2)
+The sum extends over two effective radii ($2r_{e}$), but in this case, the sum is taken within the region of the ellipse defined above.
+Here, $N$ represents the number of unmasked pixels included in the computation, $I_{i,j}$ is the galaxy intensity at pixel $(i,j)$, $S$ is the convolved surface brightness model, and $\sigma_{s}^{2}$ is the estimated photometric error, which is taken directly from the sigma image produced by GALFIT 3.
+Another useful parameter is the Tidal parameter, defined as:
+$T_{gal}=\left[\left(\frac{I_{(x,y)}}{M_{(x,y)}}\right)-1\right],$ (3)
+where $I_{(x,y)}$ are the galaxy pixel intensity values in (x, y) and $M_{(x,y)}$ represent the intensity values of the model.
+This parameter measures the differences between the galaxy image and its corresponding model.
+tal09 used this to measure the tidal disturbances in elliptical galaxies and explored their relation with environmental properties, AGN activity, and color.
+EllipSect is particularly useful to determine the total effective radius, $r_{e}^{t}$, of a galaxy when its SB model results after a multicomponent fit, for example, for a fit including a bulge, a disk, and a bar (Sérsic+Exponential+Sérsic), while GALFIT 3 only generates $r_{e}$ for each component; EllipSect calculates the total effective radius by summing all subcomponents of the galaxie’s SB model, calculating half of the total flux, and finding the corresponding radius using the bisection method.
+In the manual of GALFIT 3, the author suggests the application of a single Sérsic fit for the entire galaxy, using $r_{e}$ from the GALFIT 3 fit as a proxy for $r_{e}^{t}$.
+We argue that EllipSect significantly extends this approximation.
+To resolve this, EllipSect employs a numerical method that adds together the models of all components and then locates the radius where the total luminosity reaches half of its overall value ($r_{e}^{t}$).
+This procedure can also be applied to determine the radius corresponding to any percentage of the total light.
+EllipSect provides additional ways to determine the goodness of fit by calculating the Akaike Information Criterion (AIC, akaike74) and the Bayesian Information Criterion (BIC, schwarz78), which enable to determine the best fit for the galaxy among a given set of models.
+Both criteria penalize adding new parameters by increasing the estimate of $\chi_{\nu}$.
+The model with the lowest AIC or BIC value is considered the best fit.
+EllipSect also computes an alternative BIC using the number of resolution elements instead of the number of pixels defined by $N_{res}=\frac{n_{pix}}{\pi(\text{FWHM})^{2}}$.
+Finally, additional output parameters are computed for each model component, including the fraction of light (such as the bulge-to-total ratio for each component), $\left<\mu\right>_{e}$, $\mu_{e}$, absolute magnitude, luminosity radius at 90% of the total light, Kron radius (kron80), and Petrosian radius (petrosian76).
+Accurate determination of those parameters requires a precise estimation of the sky background.
+2.2.1 Sky computation
+An incorrect sky background estimation can lead to inaccurate surface brightness model parameters, especially the Sérsic index.
+An appropriate strategy is to determine and fix the sky background during model fitting, thereby reducing the number of free parameters and preventing parameter degeneracy.
+EllipSect provides two methods for computing the sky background: the sky gradient method and the random box method.
+GALFIT 3 does not include these two methods; therefore, by comparing the results, the user decides whether to retain the model or refine it by using the newly derived sky value from EllipSect.
+Sky gradient method: This method follows barden12.
+In this approach, EllipSect constructs concentric rings using the same angular parameters and ellipticity employed for the elliptical sectors described in 2.1.
+The user specifies the ring width and EllipSect enlarges the rings’ major axes outward from the galaxy’s center.
+EllipSect can remove the top $80\%$ and bottom $20\%$ of sky values in each ring; however, the user can also specify a mask file to exclude undesired pixels.
+Finally, EllipSect computes the gradient from the mean sky values of each ring.
+A negative gradient arises in the galaxy’s inner regions because the central counts exceed those in the outskirts.
+Beyond this zone, the gradient fluctuates randomly between positive and negative values (see Figure 5 in barden12).
+EllipSect generates a fit image with sky rings to illustrate this gradient.
+It begins calculating the gradient from an initial radius and stops when it becomes positive for the second time, following the recommendation by barden12.
+At that point, EllipSect returns the sky value determined for this radius.
+Figure 1 shows an example of an elliptical ring used to compute the sky.
+Figure 1: Example of sky computation using the gradient method.
+The ring around the brightest cluster member of Abell 2029 indicates the radius used by EllipSect to compute the sky.
+For visualization purposes, all pixels in that elliptical ring are assigned the same numerical value corresponding to the ellipse’s major-axis length.
+Random box method: For this method, EllipSect calculates the sky using randomly placed boxes around the galaxy.
+To achieve this, it first creates an elliptical mask that covers the entire galaxy.
+By default, the geometry for this elliptical mask comes from the GALFIT 3 output file, although the user may modify it.
+Next, the program positions boxes randomly between this elliptical mask and a specified maximum radius (often at the image boundary).
+It can also optionally remove the top 80% and bottom 20% of pixel values within these boxes.
+The final sky value is the mean of the median and standard deviation of all pixels within the boxes.
+The default settings place 20 boxes, each measuring 20 pixels across.
+This method follows the approach described in Gao17.
+3 Results: Examples of SB plots
+rios21 and rios25 applied EllipSect in their study of Two Micron All Sky Survey (2MASS) galaxies, where they applied GALFIT 3 to generate estimates of morphological and structural parameters through photometric decompositions.
+In another study by reyes21, EllipSect was used to map dust in the cores of local Active Galactic Nuclei (AGN), using Hubble Space Telescope (HST) images, SB models, and radiative transfer simulations.
+The code has also been used in reports, conference proceedings, professional papers, and undergraduate and graduate theses (e.g., rios21b; duggal24, Valerdi M.
+et al.
+2026, in preparation).
+Here, we present examples of SB profiles for three images: Holm 15A from the Kitt Peak National Observatory (KPNO), M61 from 2MASS, and NGC 4261 from the HST.
+3.1 Kitt Peak: Holm 15A
+This example shows the SB profile plots of the Abell 85 brightest cluster galaxy (BCG), the cD galaxy Holm 15A.
+Here, GALFIT 3 was used to fit a multi-Gaussian component to model the SB; this result complements the analysis of 2014ApJ...795L..31L, who used Nuker profiles.
+These observations were obtained with the KPNO 0.9m Telescope and are part of the LOCOS catalog (1997PhDT.......
+[^3] L).
+The reader can find details on the selection criteria, observations, and data reduction in 1997PhDT.......
+[^3] L, 2004ApJ...614..679L, and López-Cruz et al.
+(2026, in preparation).
+In Figure 2, we show the cube image generated by EllipSect for this galaxy.
+Figure 2: Example of cube image generated by EllipSect.
+The left panel shows the galaxy, the middle panel is the model, and the right panel is the residual.
+The user can change the color palette, brightness, and contrast.
+In Figure 3, we present the single-plot of the SB of Holm 15A.
+Seven Gaussians were used to cover the whole SB profile.
+The plot also shows the percent error.
+This figure was generated automatically by EllipSect, which computes the SB profiles of the galaxy, model, and individual components.
+Figure 3: Example of EllipSect output for the cD galaxy Holm 15a and its model.
+It shows a seven-Gaussian component fit.
+The red continuous line represents the galaxy and the blue continuous line represents the multicomponent GALFIT 3 model.
+Each Gaussian component is shown as a dashed line and is color-coded, as indicated in the inset box at the top right.
+The plot shows average SB vs.
+radius along the major axis.
+The residual percentage is displayed in the bottom plot.
+3.2 2MASS: M61
+In Figure 4, we show the galaxy, model, and residual for galaxy M 61 in the Ks band of 2MASS.
+Figure 4: Example of cube image generated by EllipSect for M 61 galaxy. Details for this image are the same as the caption of Figure 2.
+Figure 5 shows the multiplot of the fitting of M 61.
+This model consists of a series of Sérsic and Exponential components.
+In the region where the model fails to reproduce the SB model of the galaxy, it results in apparent deviations.
+Figure 5: Example of EllipSect multi-plot output for M 61 galaxy and its model.
+This figure shows multiple plots of galaxy SB and model at different angles from the major axis (the major axis is the one with $0\deg$).
+The error is shown on the right hand side of the multiplot.
+The angle in red at the top right indicates the angle measured relative to the image’s Y-axis.
+On the other hand, the subtle grey angle at the bottom left indicates the angle measured from the major axis, starting from the top, with the major axis and the bottom plot with the minor axis.
+3.3 HST: NGC4261
+In the last example, we display an HST image of NGC 4261. Figure 6 includes the actual galaxy, its model, and its residual.
+Figure 6: Example of cube image generated by EllipSect for NGC4261 galaxy. Details for this image are the same as the caption of Figure 2.
+Figure 7 shows the galaxy NGC4261 taken with the HST using filter F547M. This model used a single Sérsic component to fit the galaxy’s SB.
+Figure 7: Example of EllipSect output for NGC4261 galaxy and its model. The description of the image is the same as Figure 5.
+4 Discussion
+EllipSect was developed to generate surface brightness profile plots, calculate non-parametric measurements, and export data derived from GALFIT model outputs.
+Its primary purpose is to test the reliability of GALFIT 3 ’s model at various radii and angles for a given galaxy.
+In addition, EllipSect allows users to include individual components of the GALFIT 3 models in the plots, enabling verification that these components behave as expected.
+This feature streamlines the analysis process by eliminating the need to switch between different program formats, making it practical to quickly confirm that the GALFIT model is appropriate for the galaxy.
+EllipSect aids GALFIT 3 users with a readily accessible and easy-to-use tool for generating accurate SB measurements and publication-quality plots.
+One typically uses IRAF’s ellipse in the conventional approach to generating SB profiles.
+This routine fits an ellipse to each isophote at every radius, from the central region to the galaxy’s outskirts.
+As the radius changes, the ellipse’s ellipticity and position angle may also vary.
+One caveat of EllipSect is that it maintains the ellipticity and the angular position fixed throughout the entire galaxy, overlooking any possible variations of these parameters with radius.
+However, this fixed-parameter approach is advantageous when dealing with galaxies fitted simultaneously using GALFIT 3.
+In such cases, masking or removing the galaxy is insufficient, as the galaxies are blended on the line of sight, and it is challenging to distinguish where one galaxy ends, and the other begins (e.g., haussler07).
+In addition, EllipSect avoids the need for “double” fitting, i.e., two-dimensional and one-dimensional fitting, which IRAF’s ellipse requires.
+One-dimensional routines may not be capable of simultaneous fitting isophotes for two neighboring galaxies, making simultaneous fitting difficult or impossible.
+While EllipSect is not a complete substitute for IRAF’s ellipse, it offers alternative techniques to address the limitations of isophotal fitting.
+Despite the widespread use of IRAF’s ellipse, alternative techniques such as EllipSect offer unique advantages for SB analysis, which may benefit astronomers seeking to enhance their understanding of galactic structures.
+The parameter $r_{e}^{t}$ is the effective radius of the entire galaxy, which is an essential parameter derived by EllipSect; $r_{e}^{t}$ is critical to assess metallicity gradients, effective velocity dispersion ($\sigma_{eff}$) and other relevant information (e.g., 2023MNRAS.526.2479G) or comparisons with numerical models.
+However, no single component’s $r_{e}$ can accurately describe an entire galaxy, unless we deal with pure disk or bulge galaxies.
+Determining which aperture to use for a galaxy’s photometric variables can be somewhat arbitrary.
+Different photometric quantities require a suitable radius for proper evaluation, and standardizing this across various galaxies is difficult.
+In the case of GALFIT 3, the reduced chi-squared ($\chi^{2}_{\nu}$) is computed using all pixels within the fitting region.
+However, changing the size of this region can alter $\chi^{2}_{\nu}$ and may include pixels from other objects, introducing additional noise.
+Consequently, the $\chi^{2}_{\nu}$ reflects the fit quality for every object within that region.
+EllipSect uses the radius containing 90% of the total light to compute various photometric variables, $r_{9}0$ which is calculated in the same fahion as the effective radius.
+Considering the region that contains 90% of the total light minimizes the contamination from external objects.
+EllipSect calculates the Tidal parameter, Bumpiness, AIC, BIC, and a local $\chi^{2}_{\nu}$ inside $r_{9}0$.
+EllipSect provides corrections for galactic extinction and surface brightness dimming, but does not perform k-corrections at this time.
+Users can still apply morphological-based corrections to the galaxy, although this may be imprecise for cluster galaxies whose colors can differ from those of field galaxies.
+In such situations, manual correction is recommended.
+There are many ways to generate k-corrections, we found the tool http://kcor.sai.msu.ru/, which uses redshift and the color index of a given galaxy as input (chilingarian12).
+EllipSect provides a helpful useful tool for wrapping code to automate GALFIT 3 fitting process.
+It can efficiently compute the necessary photometric variables required by these codes, thereby allowing for rapid analysis of model outputs for multiple galaxies, which would otherwise be time-consuming.
+In fact, EllipSect has been also designed to be used in conjunction with the Driver for GALFIT on Cluster Galaxies (DGCG) code (Añorve et al., in preparation), which automates GALFIT 3 fitting for galaxies in dense environments.
+5 Conclusions and Future Work
+This paper introduces EllipSect algorithm implemented in Python designed to analyze outputs from GALFIT 3.
+It generates surface brightness profiles for galaxies, models, and individual components, offering single-plot and multiplot visualizations.
+Additionally, EllipSect computes various photometric parameters, including light fractions for each component, $r_{e}^{t}$, $\left<\mu\right>_{e}$, $\mu_{e}$, absolute magnitudes, the luminosity radius $90\%$, the effective radius of the galaxy, the Kron radius, and the Petrosian radius.
+It also calculates Tidal and Bumpiness parameters, signal-to-noise ratios, and local $\chi^{2}_{\nu}$ values to assess differences between models and observed galaxies.
+EllipSect offers two methods for computing the sky background and enables users to distinguish among different galaxy models by computing Bayesian information criteria.
+The code can be easily adapted to other GALFIT 3 wrapping codes and will be integrated with DGCG in future work.
+This project aims to apply EllipSect to analyze data images containing hundreds of galaxies, such as galaxy clusters.
+EllipSect is a robust tool for analyzing GALFIT 3 output and extracting extensive photometric information for galaxies.
+Its flexibility and versatility make it an essential tool for studying galaxy properties and morphology.
+Future development will focus on expanding its capacity to handle larger datasets and integrate it with other GALFIT 3 wrapping codes.
+{appendices}
+6 GALFIT 3 built-in functions or profiles
+Below, we describe the built-in functions or profiles implemented in the last distribution of GALFIT 3555See GALFIT 3 user’s manual, downloadable from https://users.obs.carnegiescience.edu/peng/work/galfit/galfit.html.
+We cite the original sources and main applications of each profile.
+2020A&A...634A.109B has shown that the Nuker profile is the most general profile, which includes the Sérsic profiles as a special case; hence, we have ordered the profiles from the more general to the particular cases.
+Our aim in including the profiles in this appendix is to homogenize the notation with that commonly used in textbooks.
   1. 1.
-    Sérsic (Sérsic, 1968): This profile is a generalization of the de Vaucouleurs profile and is typically used to fit bulges, bars, and elliptical galaxies by allowing the concentration index $n$ to vary:
+    Nuker (lauer95): This is a general profile that includes three power-law slopes, namely: $\gamma$ (inner), $\alpha$ (sharpness of the transition from inner to outer slopes), and $\beta$ (outer), they are free to vary and help to account for the large cores of very luminous early-type galaxies, including cD galaxies.
+This profile is expressed as follows:
+    $I(r)=I_{b}\,2^{(\frac{\beta-\gamma}{\alpha})}\left(\frac{r_{b}}{r}\right)^{\gamma}\left[1+\left(\frac{r}{r_{b}}\right)^{\alpha}\right]^{(\frac{\gamma-\beta}{\alpha})},$ (4)
+    where $r_{b}$ is the break radius, defined as the location at which the slope of the surface brightness profile changes, and $I_{b}$ is the SB at $r_{b}$.
+    The work of 2020A&A...634A.109B showed that the Sérsic profile is a unique subgroup of the Nuker function; hence, the de Vaucouleurs profile, the exponential profile, and the Gaussian profile discussed below become specific cases of the Nuker profile, as well.
+  2. 2.
+    Sérsic (sersic68): This profile was introduced as a generalization of the de Vaucouleurs profile and is typically used to fit bulges, bars, and elliptical galaxies by allowing the concentration index $n$ to vary:
     $I(r)=I_{e}\exp\left[-\kappa_{n}\left(\left(\frac{r}{r_{e}}\right)^{\frac{1}{n}}-1\right)\right],$ (5)
     where $I_{e}$ is the SB at the effective radius, $r_{e}$, which encloses half of the luminosity.
-The parameter $\kappa_{n}$ is related to $n$, so $r_{e}$ becomes the half-light radius; for $n\in\mathbb{R}$, $\kappa_{n}=2n-\frac{1}{3}+\frac{4}{405n}+\frac{46}{25515n^{2}}+O(n^{-3})$ gives an approximation (Ciotti & Bertin, 1999).
-See, for example, GALFIT 3 user’s manual and Graham & Driver (2005) for a full discussion on the properties of the Sérsic profile.
-  2. 2.
-    de Vaucouleurs (de Vaucouleurs, 1948): A particular case of the Sérsic profile (index $n=4$, $\kappa_{4}=-7.66925$), provides an excellent approximation to the SB profile of giant ellipticals.
+The parameter $\kappa_{n}$ is related to $n$, so $r_{e}$ becomes the half-light radius; for $n\in\mathbb{R}$, $\kappa_{n}=2n-\frac{1}{3}+\frac{4}{405n}+\frac{46}{25515n^{2}}+O(n^{-3})$ gives an approximation (1999A&A...352..447C).
+See, for example, GALFIT 3 user’s manual and graham05 for a full discussion on the properties of the Sérsic profile.
+  3. 3.
+    de Vaucouleurs (devau48): A particular case of the Sérsic profile (index $n=4$, $\kappa_{4}=-7.66925$).
+A general tendency shows that $n$ grows with galaxy luminosity; however, at high luminosities the tendency levels off ($n\rightarrow 4$) (e.g., 2003ApJ...594..186B; 2011RMxAC..39..100L; 2025arXiv250202546R).
+Thus, as originally found by devau48, this profile historically provided a widely used approximation to the SB profiles of giant ellipticals.
+However, recent studies advocate using flexible Sérsic indices (n varying) for more accurate descriptions of elliptical galaxies’ diverse structures (graham05; kormendy09).
 This profile has the following expression:
     $I(r)=I_{e}\exp\left[-7.66925\left(1-\left(\frac{r}{r_{e}}\right)^{1/4}\right)\right]$ (6)
-  3. 3.
-    Exponential (Patterson, 1940): This popular profile is used to fit the disks of spiral galaxies; it’s a particular case of the Sérsic profile with $n=1$:
+  4. 4.
+    Exponential (1940BHarO.914....9P): This popular profile is used to fit the disks of spiral galaxies; it’s a particular case of the Sérsic profile with $n=1$:
     $I(r)=I_{0}\exp\left(-\frac{r}{r_{s}}\right),$ (7)
-    where $I_{0}$ is the central SB and $r_{s}$ is called the scale length, which is related to the effective radius by $r_{e}=1.67835\,r_{s}$ for $n=1$.
-Freeman (1970) proposed a combination of a de Vaucouleurs bulge and an exponential disk to fit late-type galaxies.
-  4. 4.
-    Nuker (Lauer et al., 1995): This is a general profile that includes three power-law slopes, namely: $\gamma$ (inner), $\alpha$ (sharpness of the transition from inner to outer slopes), and $\beta$ (outer), they are free to vary and help to account for the large cores of very luminous early-type galaxies, including cD galaxies.
-This profile is expressed as follows:
-[…]
-    where $r_{b}$ is called the break radius, and it’s the place where the curvature of the profile reaches its maximum and $I_{b}$ is the SB at $r_{b}$.
-    Recently, Baes (2020) showed that the Sérsic profile is a unique subgroup of the Nuker function; hence, the de Vaucouleurs profile, the exponential profile, and the Gaussian profile discussed above become specific cases of the Nuker profile, as well.
+    $I_{0}$ is the central SB and $r_{s}$ is called the scale length, which is related to the effective radius by $r_{e}=1.67835\,r_{s}$ for $n=1$.
+1970ApJ...160..811F proposed a combination of a de Vaucouleurs bulge and an exponential disk to fit late-type galaxies.
   5. 5.
-    Edge-on disk (van der Kruit & Searle, 1981): This profile is used to fit the vertical structure of edge-on spiral galaxies.
-It results after considering an locally isothermal thin disk (van der Kruit & Searle, 1981):
+    Gaussian profile is used to fit PSF profiles, the contribution of active galactic nuclei and bars; it’s a special case of the Sérsic profile when $n=0.5$. The following equation describes it:
+    $I(r)=I_{0}\exp\left(\frac{-r^{2}}{2\sigma^{2}}\right),$ (8)
+    The full width parametrizes the scale at half maximum ${\rm FWHM}=2\sqrt{2\ln{2}}\approx 2.355\,\sigma$, instead of $r_{e}$.
+  6. 6.
+    Edge-on disk (1981A&A....95..105V): This profile fits the vertical structure of edge-on spiral galaxies. It results after considering an locally isothermal thin disk (1981A&A....95..105V):
     $I(r,z)=I_{0}\left(\frac{r}{r_{s}}\right)K_{1}\left(\frac{r}{r_{s}}\right)\operatorname{sech}^{2}\left(\frac{z}{z_{h}}\right),$ (9)
-    where $I_{0}=I(0,0)$ is the central SB, $r_{s}$ is the scale length on the disk’s major axis, $z_{s}$ is the scale high on the $z$ axis perpendicular to the disk, and $K_{1}$ is the modified Bessel function of the second kind.
-  6. 6.
-    Ferrers [^4] Peng et al.
-(2010) called it modified Ferrer profile.: This model is named after the triaxial potential introduced by Norman Macleod Ferrers (Ferrers, 1877; Binney & Tremaine, 1987, pp.
+    where $I_{0}=I(0,0)$ is the central SB, $r_{s}$ is the scale length on the disk’s major axis, $z_{s}$ is the scale height on the $z$ axis perpendicular to the disk, and $K_{1}$ is the modified Bessel function of the second kind.
+  7. 7.
+    Ferrers666peng10 called it modified Ferrer profile.: This model is named after the triaxial potential introduced by Norman Macleod Ferrers (Ferrers1877; 1987gady.book.....B, pp.
 61-62).
@@
 Notwithstanding, a Sérsic profile whose index is $n<0.5$ approximates the Ferrers profile; therefore, either profile can be used to model the SB of bars.
-Besides, the results are indistinguishable (e.g., Kim et al., 2015).
-  7. 7.
-    Empirical King (Elson, 1999): It is given as follows:
-    $I(r)=I_{0}\left[1-\frac{1}{\left(1-\left(\frac{r_{t}}{r_{c}}\right)^{2}\right)^{\frac{1}{\alpha}}}\right]^{-\alpha}\left[\frac{1}{\left(1-\left(\frac{r}{r_{c}}\right)^{2}\right)^{\frac{1}{\alpha}}}-\frac{1}{\left(1-\left(\frac{r_{t}}{r_{c}}\right)^{2}\right)^{\frac{1}{\alpha}}}\right]^{\alpha},$ (7)
-    where the core radius $r_{c}$, the truncation radius $r_{t}$, and the power law index $\alpha$ are free parameters; however, in the standard modified King profile, $\alpha=2$ is fixed.
+Besides, the results are indistinguishable (e.g., 2015ApJ...799...99K).
+  8. 8.
+    Empirical King (1999glcl.conf..209E): It is given as follows:
+    $I(r)=I_{0}\left(1-\frac{1}{\psi^{\frac{1}{\alpha}}}\right)^{-\alpha}\left(\frac{1}{\xi^{\frac{1}{\alpha}}}-\frac{1}{\psi^{\frac{1}{\alpha}}}\right)^{\alpha};$ (11)
+    where $\psi=\left(1-\frac{r_{t}}{r_{c}}\right)^{2}$ and $\xi=\left(1-\frac{r}{r_{c}}\right)^{2}$, while the core radius $r_{c}$, the truncation radius $r_{t}$, and the power law index $\alpha$ are free parameters; however, in the standard modified King profile, $\alpha=2$.
 This profile fits the SB of globular clusters.
-  8. 8.
-    Gaussian profile is used to fit PSF profiles, the contribution of active galactic nuclei and bars; it’s a special case of the Sérsic profile when $n=0.5$.
-It is described by the following equation:
-    $I(r)=I_{0}\exp\left(\frac{-r^{2}}{2\sigma^{2}}\right),$ (8)
-    the scale is parametrized by the full with at half maximum ${\rm FWHM}=2\sqrt{2\ln{2}}\approx 2.355\,\sigma$, instead of $r_{e}$.
   9. 9.
-    Moffat (Moffat, 1969) is a profile used to fit PSF profiles or the emission of active nuclei of galaxies. The following equation describes it:
+    Moffat (1969A&A.....3..455M) is a profile used to fit PSF profiles or the emission of active nuclei of galaxies. The following equation describes it:
     $I(r)=\frac{I_{0}}{\left(1+\left(\frac{r}{r_{d}}\right)^{2}\right)^{n}},$ (12)
-    where $n$ is the concentration index; as for the Gaussian profile, the scale is given by $FWHM=2\left(2^{\frac{1}{n}}-1\right)^{\frac{1}{2}}\,r_{d}$.
+    where nn is the concentration index; as for the Gaussian profile, the scale is given by $FWHM=2\left(2^{\frac{1}{n}}-1\right)^{\frac{1}{2}}\,r_{d}$.
 Nevertheless, the Gaussian profile is more widely used than the Moffat.
@@
   12. 12.
-    Spiral modes are used to model the spiral arms or asymmetrical galaxies; see Peng et al. (2010) and GALFIT 3 user’s manual for more details.
-With GALFIT 3, the user can apply multiple profiles to fit the following components: bulges, disks, bars, spiral arms, and asymmetries to one or several galaxies simultaneously.
-This gives a substantial advantage in crowded galaxy fields, such as compact groups or clusters of galaxies.
-Galaxies being extended objects are modeled better when the galaxy of interest is fitted together with its closely neighboring galaxies instead of masking.
-This has been shown through simulations (Häussler et al., 2007; Añorve, 2012).
-GALFIT 3 applies a $\chi^{2}_{\nu}$ optimization scheme for parameter selection, including the background, generating parameter uncertainties, and produces a FITS (Flexible Image Transport System) image cube containing the image of the input object, the model image resulting after the fit, and the residual image.
-However, GALFIT 3 does not include a graphical tool to compare the input image with its resulting model image.
-Users commonly applied IRAF’s ellipse; nevertheless, this suite of programs became difficult to access after IRAF’s discontinuation in 2013; fortunately, IRAF’s ellipse is implemented in the Astropy library and new releases through.
-In any case, the main difficulty is that one needs to adequate GALFIT 3 ’s output format to the format admitted by external codes.
-However, this process could be time-consuming and prone to mistakes in handling.
-Additionally, in crowded fields, neighboring galaxies near the target galaxy can be fitted simultaneously using GALFIT 3, resulting in more accurate outcomes.
-Häussler et al.
-(2007) demonstrated through simulations that masking galaxies in such fields leads to inaccurate fits for both the target galaxy and its neighbors.
-Improved fits were achieved when galaxies nearby were fitted simultaneously.
-However, this method is not feasible with IRAF’s ellipse task, which processes only one galaxy at a time.
-Hence, we are introducing EllipSect, a new Python tool that generates SB profiles and extracts complementary information directly from GALFIT 3 ’s output.
-We provide an easy-to-use tool to help users select and improve their models by introducing additional components or removing redundant components.
-EllipSect ’s output includes data from the observed galaxy, and its model fits in the form of one-dimensional profiles.
-For a more straightforward assessment of the results, EllipSect can handle individual model components for a detailed analysis.
-Additionally, EllipSect complements GALFIT 3 analysis by calculating integrated magnitudes, mean SB at $r_{e}$ ($\mu\right>_{e}$), the SB at $r_{e}$ ($\mu_{e}$), the radius at 90% of total light, Kron radius (Kron, 1980), Petrosian radius (Petrosian, 1976), bulge-to-total ratio, and the total effective radius ($r_{e}^{t}$) of a galaxy when multi-components are used.
-This paper is organized as follows: §1 presents a general introduction to GALFIT 3 and the built-in profiles, §2 explains EllipSect ’s implemented algorithms.
-§3 shows applications and examples, §4 and §5 presents a discussion and conclusions, respectively.
-The Appendices show how to install and run EllipSect.
-0.2 EllipSect
-This section explains how EllipSect works.
-We have divided it into two parts: first, we explain the main algorithm, and then, we explain the photometric measurements.
-Throughout this article, we refer to the individual profiles that make a SB model the components, while the set of SB components is called the model.
-0.2.1 The Algorithm
-EllipSect was designed to be user-friendly for most users.
-It omits any direct interaction with the code or translation of GALFIT 3 ’s data format.
-The program is efficiently executed via the command line.
-It only requires the latest GALFIT 3 ’s model output files.
-GALFIT 3’s output files consist of a data cube in FITS format, which includes the original image but chopped into the fit region, the model image, and residual image (i.e., galaxy image -- model image), a file named galfit.nn (where nn indicates the number of attempts run by GALFIT 3).
-This last file contains the parameters that fit each component of the model.
-Additionally, GALFIT 3 creates a log file including the model parameter values and their errors.
-EllipSect uses the GALFIT output file (galfit.nn) to parse the input magnitude zero point, plate scale, mask file name, output file, and model parameters.
-The output GALFIT 3 file may contain information about the simultaneous fit of neighboring galaxies to the target galaxy.
-Using the distances among model components, EllipSect can discern which components belong to each galaxy.
-EllipSect uses the same mask employed by GALFIT to remove the undesired pixels from the image.
-Once this is done, EllipSect reads the sky component from the GALFIT output file and subtracts it from the output galaxy and model images.
-The program divides the galaxy and its model into sectors equally spaced in angle around their centers.
-This is accomplished using the sectors_photometry routine from the mgefit555mgefit is a software to obtain Multi-Gaussian Expansion (MGE) fits galaxy images (Cappellari, 2002) Python library by Cappellari (2002).
-The routine also calculates the counts within each sector, extending outward testing the galaxy’s center until a specified minimum count level is reached.
-This is set by the MINLEVEL parameter of the routine.
-The default value is zero for this parameter.
-The ellipse is divided into sectors from the major axis to the minor axis, with the number of sectors set to a default value of $19$, resulting in sectors spanning an angle of $5^{\circ}$.
-Sectors around the galaxy are built using the position angle and ellipticity found in the model component of the output GALFIT file.
-If multiple components are fitted to the galaxy, the geometric information is taken from the last component in the model, as it is identified as the galaxy’s outermost component.
-This component is typically an exponential model, although this may vary depending on the specific galaxy model used.
-The user can change all these default parameters.
-EllipSect generates publication-quality graphs.
-It produces two basic graphs in its simplest mode: one showing the averaged surface brightness along the major axis and another displaying the SB for various azimuthal angles across multiple plots (see examples in the Results section).
-Additionally, this mode generates a png image featuring the galaxy, model, and residual.
-EllipSect can also export a file containing the data points for each plot, allowing users to create custom visualizations with other plotting routines.
-Optionally, the SB profile for each plot component can include the galaxy (via the $--comp$ flag).
-This feature allows users to visualize the behavior of model components and identify irrelevant components or redundancies (see, for example, Figure 3).
-Based on the mgefit manual, the SB is calculated according to the following formula666this does not include the extinction parameter :
-$\mu_{I}=zpt+5\times\log(Pl_{sce})+2.5\times\log(t_{exp})-2.5\times\log(C_{0})+0.1,$ (10)
-where $zpt$ is the magnitude zero point, $C_{0}$ are the $counts\times pixel^{-1}$, $Pl_{sce}$ is the plate scale, $t_{exp}$ is the exposition time, and $0.1$ is a correction for infinite aperture.
-0.2.2 Output photometry
-In addition to generating surface brightness plots, EllipSect calculates additional photometric variables from the models beyond those provided by GALFIT 3.
-These variables are stored in an output file and are derived after EllipSect extracts the count values for each sector.
-This section describes these variables and explains their computation.
-EllipSect connects to NED through the command curl to download a xml file to extract the photometric corrections and compute the absolute magnitude according to the appropriate band.
-In the case of luminosity, it is calculated using the absolute magnitude of the Sun taken from Willmer (2018).
-This file extracts the distance modulus and the galactic extinction.
-It also obtains a conversion parameter used to compute the radius in Kpc units from angular units.
-SB dimming also corrects the mean SB at effective radius $\mu\right>_{e}$ and SB at the effective radius $\mu_{e}$.
-The only correction not applied is the k-correction since it needs the color information that is not always available 777see http://kcor.sai.msu.ru/.
-However, correction through galactic morphology can be used as well.
-It is important to note that corrections are applied only to the quantities in the output file; the single and multiple plots mentioned earlier are not adjusted for these corrections.
-However, users can apply a correction factor as a command-line argument or generate plots using the data from the EllipSect output files.
-NED provides two distance modulus measurements: one derived from cosmological models and another obtained directly from the galaxy.
-By default, EllipSect uses the redshift-independent measurement to calculate the luminosity, if available.
-Otherwise, it uses the distance modulus based on cosmological models.
-Most output photometric variables are calculated within the ellipse corresponding to the radius enclosing $90\%$ of the total light.
-Key variables include Bumpiness (Blakeslee et al., 2006), tidal parameter (Tal et al., 2009), reduced chi-squared ($\chi^{2}_{\nu}$) within this ellipse (as opposed to the value provided by GALFIT 3, which applies to the entire fitted image section), signal-to-noise ratio (S/N), residual sum of squares, number of degrees of freedom, number of free parameters, absolute magnitude, the bulge-to-total ratio, and luminosity.
-The Bumpiness parameter quantifies the irregular structure of a galaxy by measuring deviations in its SB from a smooth SB model.
-Blakeslee et al.
-(2006) employed this parameter and the Sérsic index to classify galaxies into elliptical, S0, spiral, and irregular types.
-The Bumpiness parameter is computed as follows:
-$B=\frac{\left(\frac{1}{N}\sum_{r_{i,j}<2r_{e}}[I_{i,j}-S(x_{i},y_{j}|r_{e},n)]_{s}^{2}-\sigma_{s}^{2}\right)^{\frac{1}{2}}}{\frac{1}{N}\sum_{r_{i,j}<2r_{e}}S(x_{i},y_{j}|r_{e},n)},$ (11)
-The sum is performed over two effective radii ($2r_{e}$), but in this case, it is applied within the ellipse region defined above.
-Here, $N$ represents the number of unmasked pixels included in the computation, $I_{i,j}$ is the galaxy intensity at pixel $(i,j)$, $S$ is the convolved surface brightness (SB) model, and $\sigma_{s}^{2}$ is the estimated photometric error, which is taken directly from the sigma image produced by GALFIT 3.
-Another useful parameter is the Tidal (Tal et al., 2009) parameter is defined as:
-$T_{gal}=\left[\left(\frac{I_{(x,y)}}{M_{(x,y)}}\right)-1\right],$ (12)
-where $I_{(x,y)}$ are the galaxy pixel intensity values in (x, y) and $M_{(x,y)}$ represent the intensity values of the model.
-This parameter measures the differences between the galaxy image and its corresponding model.
-Tal et al.
-(2009) used this to measure the tidal disturbances in elliptical galaxies and explored its relation with environmental properties, AGN activity, and color.
-Additionally, suppose the galaxy has been fitted with multiple Sérsic components (or variants such as Gaussian and exponential models).
-In that case, the program gives $r_{e}^{t}$.
-While GALFIT 3 provides $r_{e}$ for each component, the program determines $r_{e}^{t}$ by summing the surface brightness model of all components, calculating half of the total flux, and identifying the corresponding radius using the bisection method.
-The Akaike Information Criterion (AIC) (Akaike, 1974) and the Bayesian Information Criterion (BIC) (Schwarz, 1978) determines which model best fits the galaxy among a given set.
-Both criteria penalize adding new parameters by increasing the estimate of $\chi_{\nu}$.
-The model with the lowest AIC or BIC value is considered the best fit.
-EllipSect computes both criteria and also calculates an alternative Bayesian Information Criterion (BIC) using the number of resolution elements instead of the number of pixels defined by $N_{res}=\frac{n_{pix}}{\pi(\text{FWHM})^{2}}$.
-Finally, additional output parameters are computed for each model component, including the fraction of light (such as the bulge-to-total ratio for each component), $\mu\right>_{e}$, $\mu_{e}$, absolute magnitude, luminosity radius at 90% of the total light, Kron radius (Kron, 1980), and Petrosian radius (Petrosian, 1976).
-Sky computation
-An incorrect sky background estimation can lead to inaccurate surface brightness model parameters, especially the Sérsic index.
-An appropriate strategy is determining the sky background and fixing it during model fitting.
-This reduces the number of free parameters and prevents parameter degeneracy.
-EllipSect provides two methods for computing the sky background: The Sky gradient method and the random box method.
-These methods are not used to estimate the photometric variables discussed above.
-Instead, they enable a comparison with the sky value computed by GALFIT 3, allowing the user to decide whether to refine the model using the newly derived sky value from EllipSect.
-Sky gradient method.
-This method follows Barden et al.
-(2012).
-In this approach, EllipSect constructs concentric rings using the same angular parameters and ellipticity employed for the elliptical sectors described in 0.2.1.
-The user specifies the ring width, and EllipSect enlarges the rings’ major axes outward from the galaxy’s center.
-A mask file is used to exclude undesired pixels.
-Optionally, EllipSect can remove the top $80\%$ and bottom $20\%$ of sky values in each ring.
-Finally, it computes the gradient based on the mean sky values of each ring.
-A negative gradient arises in the galaxy’s inner regions because the central counts exceed those in the outskirts.
-Beyond this zone, the gradient fluctuates randomly between positive and negative values (see Figure 5 in Barden et al.
-(2012)).
-EllipSect generates a fit image with sky rings to illustrate this gradient.
-It begins calculating the gradient from an initial radius and stops when it becomes positive for the second time, following the recommendation by Barden et al.
-(2012).
-At that point, EllipSect returns the sky value determined for this radius.
-Figure 1 shows an example of an elliptical ring used to compute the sky.
-Figure 1: Example of sky computation using the gradient method.
-The ring around the bright cluster member of Abell 2029 indicates the radius at which the sky was computed for this galaxy.
-For visualization purposes, all pixels in that elliptical ring are assigned the same numerical value corresponding to the ellipse’s major-axis length.
-Random box method
-For this method, EllipSect calculates the sky using randomly placed boxes around the galaxy.
-To achieve this, it first creates an elliptical mask that covers the entire galaxy and excludes pixels in the original input mask file.
-By default, the geometry for this elliptical mask comes from the GALFIT 3 output file, although the user may modify it.
-Next, the program positions boxes randomly between this elliptical mask and a specified maximum radius (often at the image boundary).
-It can also optionally remove the top 80% and bottom 20% of pixel values in these boxes.
-The final sky value is derived by computing the mean, median, and standard deviation of all pixels in the boxes.
-The default settings place 20 boxes, each measuring 20 pixels across.
-This method follows the approach described in Gao & Ho (2017).
-0.3 Results: Examples of SB plots
-EllipSect has been used to analyze galaxy images from 2MASS (Two Micron All Sky Survey) in which estimations of morphological and structural parameters have been obtained through photometric decompositions using GALFIT 3 (Ríos-López et al., 2021).
-Moreover, EllipSect was used in a study of dust nuclear structures of a sample of Active Galactic Nuclei (AGN) in the local Universe through the Hubble Space Telescope (HST) images, SB models, and radiative transfer simulations (Reyes-Amador, 2021).
-The code has also been used in reports, conference proceedings, professional papers, and undergraduate and graduate theses (Ríos-López, 2021, Valerdi M.
-et al.
-paper in progress; Ríos-López et al.
-paper in progress).
-Here, we present examples of surface brightness profile plots for three images: Holm 15A with Kitt Peak National Observatory (KPNO), M61 with 2MASS, and NGC4261 with the Hubble Space Telescope (HST).
-0.3.1 Kitt Peak: Holm 15A
-This example shows the SB profile plots of Abell 85 brightest cluster galaxy (BCG), the cD galaxy Holm 15A.
-Here, GALFIT 3 was used to fit a multi-gaussian component to model its SB; another analysis using the Nuker profiles has already been published (López-Cruz et al., 2014).
-This frame was taken with KPNO 0.9m Telescope, part of the catalog of LOCOS (López-Cruz, 1997).
-Details on the selection criteria, observations, and data reduction can be found in López-Cruz (1997), López-Cruz et al.
-(2004), and López-Cruz et al.
-(2025, in preparation).
-In Figure 2, we show the cube image generated by EllipSect for this galaxy.
-Figure 2: Example of cube image generated by EllipSect.
-The left panel shows the galaxy, the middle panel is the model, and the right is the residual.
-The user can change the color palette, brightness, and contrast.
-In Figure 3, we present the single-plot and multiplot of the surface brightness for the brightest cluster member of Abell 85.
-This galaxy was modeled using seven Gaussians to capture its full SB profile.
-The multiplot displays the error, while the single-plot shows the percent error.
-These plots were generated automatically by EllipSect, which computes the SB profiles of the galaxy, model, and individual components similar to IRAF’s ellipse task.
-Figure 3: Example of EllipSect output for the elliptical galaxy Holm 15a and its model.
-This was fitted using seven Gaussian components.
-In both panels, red represents the galaxy and blue represents the GALFIT 3 model.
-The color for each Gaussian component is shown in the top right inset box.
-These colors represent the same for both panels.
-Left panel: average SB vs.
-radius along the major axis.
-The residual percentage is shown in the bottom left plot.
-The model errors depend on the average over the ellipse for the model in that radius.
-Right panel: Multiple plots of galaxy SB and model at different angles from the major axis (the major axis is the one with $0\deg$).
-The error is shown on the right side of the multiplot.
-The angle in red at the top right indicates the angle measured from the Y-axis of the image.
-On the other hand, the subtle grey angle at the bottom left indicates the angle measured from the major axis starting from the top with the major axis and the bottom plot with the minor axis
-0.3.2 2MASS: M61
-In Figure 4, we show the galaxy, model, and residual for galaxy M 61 in the Ks band of 2MASS.
-Figure 4: Example of cube image generated by EllipSect for M 61 galaxy. Details for this image are the same as the caption of Figure 2.
-Figure 5 shows the multiplot of the fitting of M 61.
-This model consists of a series of Sérsic and Exponential components.
-As can be seen, deviations from the model can be appreciated in the region where the model fails to reproduce the SB model of the galaxy.
-Figure 5: Example of EllipSect output for M 61 galaxy and its model. The description of the image is the same as in the right panel of Figure 3
-0.3.3 HST: NGC4261
-In the last example, we show an image from HST; this is NGC4261 and in Figure 6 is shown the galaxy, model, and residual.
-Figure 6: Example of cube image generated by EllipSect for NGC4261 galaxy. Details for this image are the same as the caption of Figure 2.
-Figure 7 shows the galaxy NGC4261 taken with the HST using filter F547M. This model used a single Sersic component to fit the galaxy’s SB.
-Figure 7: Example of EllipSect output for NGC4261 galaxy and its model. The description of the image is the same as the right panel of Figure 3.
-0.4 Discussion
-EllipSect was developed as an extension to GALFIT 3 to generate SB profile plots and offer additional photometric information beyond GALFIT 3.
-Its primary purpose is to test the reliability of GALFIT 3 ’s model at various radii and angles for a given galaxy.
-In addition, EllipSect allows users to include individual components of the GALFIT 3 models in the plots, enabling verification that these components behave as expected.
-This feature streamlines the analysis process by eliminating the need to switch between different program formats, making it practical to quickly confirm that the GALFIT model is appropriate for the galaxy.
-EllipSect aids GALFIT 3 users with a readily accessible and easy-to-use tool for generating accurate SB measurements and publication-quality plots.
-Using the sectors_photometry routine, EllipSect creates evenly spaced angular sectors to compute surface brightness at various radii.
-It divides an ellipse, centered on the galaxy, into multiple sectors according to the geometric information from the GALFIT 3 output file.
-By default, EllipSect derives the angular orientation and ellipticity from the last component of the GALFIT 3 output file (see 0.6.2).
-In the conventional approach to generating SB profiles, one typically uses IRAF’s ellipse.
-This routine fits an ellipse to each isophote at every radius, from the central region to the galaxy’s outskirts.
-As the radius changes, the ellipticity and position angle of the ellipse may also vary.
-One disadvantage of EllipSect is that it maintains a fixed ellipticity and angular position values throughout the entire galaxy, which may hinder studying how these parameters change with radius.
-However, this fixed parameter approach is advantageous when dealing with galaxies simultaneously fitted by GALFIT 3.
-In such cases, masking or removing the galaxy is insufficient, as the galaxies are blended in the line of sight, and it is challenging to distinguish where one galaxy ends and the other begins (Haussler et al.
-2007).
-In addition, EllipSect avoids the need for ”double” fitting, i.e., two-dimensional and one-dimensional fitting, which IRAF’s ellipse requires.
-One-dimensional routines may not be capable of simultaneous fitting of isophotes for two neighboring galaxies, making simultaneous fitting difficult or impossible.
-While EllipSect is not a complete substitute for IRAF’s ellipse, it offers distinct advantages depending on the scientific question being addressed.
-Despite the widespread use of IRAF’s ellipse, alternative techniques such as EllipSect offer unique advantages for SB analysis, which may prove to be beneficial for astronomers seeking to enhance their understanding of galactic structures.
-An important output parameter from EllipSect is $r_{e}^{t}$, the effective radius of the entire galaxy, which can be used to parameterize metallicity gradients and other relevant information (e.g., Garduño et al., 2023).
-If the galaxy has only a single Sérsic component, $r_{e}$ from that component can represent the whole galaxy.
-However, no single component’s $r_{e}$ can accurately describe the entire galaxy when multiple components are present, because galaxies do not perfectly follow the surface models.
-To resolve this, EllipSect employs a numerical method that adds together the models of all components and then locates the radius where the total luminosity reaches half of its overall value ($r_{e}^{t}$).
-This procedure can also be applied to determine the radius corresponding to any percentage of the total light.
-Determining which aperture to use for a galaxy’s photometric variables can be somewhat arbitrary.
-Different photometric quantities require a suitable radius for proper evaluation, and standardizing this across various galaxies is difficult.
-In the case of GALFIT 3, the reduced chi-squared ($\chi^{2}_{\nu}$) is computed using all pixels within the fitting region.
-However, changing the size of this region can alter $\chi^{2}_{\nu}$ and may include pixels from other objects, introducing additional noise.
-Consequently, the $\chi^{2}_{\nu}$ reflects the fit quality for every object within that region.
-In contrast, EllipSect uses the radius containing 90% of the total light to compute various photometric variables.
-This radius is found by applying the same procedure used to determine the effective radius but at 90% of the total light.
-Focusing on this region ensures that only pixels belonging to the galaxy are included, minimizing contamination from external objects.
-Within this radius, EllipSect calculates several photometric variables, such as the tidal parameter, Bumpiness, AIC, BIC, and a local $\chi^{2}_{\nu}$.
-EllipSect provides corrections for galaxy extinction and surface brightness (SB) dimming, but does not perform k-corrections at this time.
-Users can still apply morphological-based corrections to the galaxy, although this may be imprecise for cluster galaxies whose colors can differ from those of field galaxies.
-In such situations, manual correction is recommended.
-Note that k-corrections can be generated using online tools such as http://kcor.sai.msu.ru/.
-EllipSect provides a helpful tool for wrapping codes aimed at automating the GALFIT 3 fitting process.
-It can efficiently compute the necessary photometric variables required by these codes, thereby allowing for rapid analysis of model outputs for multiple galaxies, which would otherwise be time-consuming.
-In fact, EllipSect has been specifically designed to be used in conjunction with the Driver for GALFIT on Cluster Galaxies (DGCG) code (Añorve et al., in preparation), which automates GALFIT 3 fitting for galaxies in cluster images.
-0.5 Conclusions and Future Work
-The paper introduces EllipSect, a Python tool for analyzing GALFIT 3 output files.
-It generates surface brightness profiles for galaxies, models, and individual components, offering single-plot and multiplot visualizations.
-Additionally, EllipSect computes various photometric parameters, including light fractions for each component, $r_{e}^{t}$, $\mu\right>_{e}$, $\mu_{e}$, absolute magnitudes, the luminosity radius $90\%$, the effective radius of the galaxy, the Kron radius, and the Petrosian radius.
-It also calculates tidal and Bumpiness parameters, signal-to-noise ratios, and local $\chi^{2}_{\nu}$ values to assess differences between models and observed galaxies.
-EllipSect offers two methods for computing the sky background and enables users to distinguish among different models for galaxies by calculating the Akaike and Bayesian information criteria.
-The code can be easily adapted to other GALFIT 3 wrapping codes and will be integrated with DGCG in future work.
-This project aims to apply EllipSect to analyze data images containing hundreds of galaxies, such as galaxy clusters.
-EllipSect is a robust tool for analyzing GALFIT 3 output and extracting extensive photometric information for galaxies.
-Its flexibility and versatility make it an essential tool for studying galaxy properties and morphology.
-Future development will focus on expanding its capacity to handle larger datasets and integrate it with other GALFIT 3 wrapping codes.
-{appendices}
-0.6 How to install and run EllipSect
-0.6.1 Installing EllipSect
-The code is written for Python 3. It can be downloaded from the site:
-https://github.com/canorve/EllipSect
-Quick installation
+    Spiral modes are used to model the spiral arms or asymmetrical galaxies; see peng10 and GALFIT 3 user’s manual for more details.
+7 How to install and run EllipSect
+7.1 Installing EllipSect
+EllipSect is implemented in Python 3 for the analysis of GALFIT’s outputs and the generation of scientific figures, and is available at https://github.com/canorve/EllipSect
+7.1.1 Quick installation
 The quickest way to install EllipSect is just by typing this into the command line:
-1pip install ellipse
-Step-by-step installation
-The following libraries are required to install EllipSect (v3.4.4) along with Python >=>=> = v3.8:
-  • numpy >=>=> = v1.20.3
-  • astropy >=>=> = v5.1
-  • scipy >=>=> = v1.5.2
-  • matplotlib >=>=> = v3.5.2
-  • mgefit >=>=> = v5.0.13
-  • sphinx >=>=> = v3.2.1
+pip install ellipse
+7.1.2 Step-by-step installation
+The following libraries are required to install EllipSect (v3.4.4) along with Python >=>= v3.8 (python2009):
+  • numpy >=>= v1.20.3 (Numpy2020)
+  • astropy >=>= v5.1 (astropy:2013; astropy:2018)
+  • scipy >=>= v1.5.2 (Scipy2020)
+  • matplotlib >=>= v3.5.2 (matplotlib2007)
+  • mgefit >=>= v5.0.13 (cappellari02)
+  • sphinx >=>= v3.2.1 (sphinx2021)
 After installing the libraries, download the latest release from the link shown above, and install it via:
-1cd EllipSect
-2pip install .
+cd EllipSect
+pip install .
 or
-1cd EllipSect
-2python setup.py install
+cd EllipSect
+python setup.py install
 To check that everything is working properly, run the automated tests:
-1python -m pytest
-Although GALFIT is not strictly required, it will require to create the model components and sigma image. Ensure you can call GALFIT from the command line. Otherwise, the automated tests will fail.
-0.6.2 Running EllipSect
+python -m pytest
+Although GALFIT is not strictly required, it will require to create the model components and sigma image. Ensure you can call GALFIT from the command prompt. Otherwise, the automated tests will fail.
+7.2 Running EllipSect
 Easy run: Once installed, run the command EllipSect in the same directory that you run GALFIT. The easiest way to run the program is:
-1EllipSect GALFIT.01
-Note: EllipSect takes the GALFIT output files (GALFIT.XX) where XX is the output file number you want to plot.
+EllipSect GALFIT.01
+Note: EllipSect takes the GALFIT output files (GALFIT.nn) where nn is the output file number you want to plot.
 More options: As explained in the paper, EllipSect uses an ellipsed-based grid to compute the SB.
 The information of this ellipse is obtained from the GALFIT output file, but if the user wants to provide their ellipse parameters, this is done via:
-1EllipSect GALFIT.01 -q 0.5 -pa 80
+EllipSect GALFIT.01 -q 0.5 -pa 80
 In this last example, an ellipse with a $0.5$ axis ratio and position angular of $80$ (measured from Y-axis) was used to create the plots.
 In case your model contains two or more components, the comp (or cp) option will plot each component (for an example of this check figure 3):
-1EllipSect GALFIT.01 -cp
+EllipSect GALFIT.01 -cp
 There are many options to change the format of the plots such as: logx, ranx, rany, pix, grid, dpi, galax, but if the user prefers to make their plots the option sbout will create a file (or files) containing the points of the the SB of the galaxy and model at each radius.
 To compute the photometry of the galaxy and model such as snr, BIC, AIC, absolute Magnitude, Luminosity, Bumpiness, Tidal and other photometric variables the option phot creates a file with the computed variables:
-1 EllipSect GALFIT.01 --phot
-EllipSect has the option to compute the sky to compare with or provide to GALFIT.
+EllipSect GALFIT.01 phot
+EllipSect has the option to compute the sky level to compare with or to provide to GALFIT.
 As explained above, the program uses two methods to compute it: The gradient method and the random box method.
 To compute the gradient method:
-1 EllipSect GALFIT.01 -gsky
+EllipSect GALFIT.01 -gsky
 And to compute using the random box method:
-1 EllipSect GALFIT.01 -rsky
+EllipSect GALFIT.01 -rsky
 The ring’s width and box’s size can be modified from their default options.
 For more options and to see how to use them, type:
-1EllipSect --help
-Running EllipSect as a script
+EllipSect help
+7.2.1 Running EllipSect as a script
 If you want to use EllipSect inside your Python script, you can call it in the following way:
-1 from ellipsect.inout.read import ArgParsing
-2 from ellipsect.sectors.sect import SectorsGalfit
-3
-4 args=[’GALFIT.01’,’--logx’, ’--phot’,’--noplot’]
-5 parser_args = ArgParsing(args)
-6 photapi = SectorsGalfit(parser_args)
-7
-8 print("Akaike Criterion: ", photapi.AICrit)
-9 print("Bulge to Total: ", photapi.BulgeToTotal)
 Listing 1: Example Python Code with EllipSect
+from ellipsect.inout.read import ArgParsing
+from ellipsect.sectors.sect import SectorsGalfit
+args=[’GALFIT.01’,’–logx’, ’–phot’,’–noplot’]
+parser_args = ArgParsing(args)
+photapi = SectorsGalfit(parser_args)
+print(”Akaike Criterion: ”, photapi.AICrit)
+print(”Bulge to Total: ”, photapi.BulgeToTotal)
 A manual and more details can be found in https://github.com/canorve/EllipSect